### v2

# SymCERE: Symmetric Contrastive Learning for Robust Review-Enhanced Recommendation

Abstract.
Modern recommendation systems can achieve high performance by fusing user behavior graphs (via GNNs) and review texts (via LLMs). However, this fusion faces three significant issues: (1) False Negatives in contrastive learning can degrade the training signal by penalizing similar items; (2) Popularity Bias, often encoded as embedding magnitude, can distort similarity scores; and (3) Signal Ambiguity, which arises from the conflation of objective facts with subjective sentiment in reviews. These interconnected issues can prevent models from learning users’ true preferences.
In this paper, we propose SymCERE (Symmetric SINCERE), a contrastive learning method that addresses these three issues simultaneously through its structural design. First, we introduce a symmetric application of the SINCERE loss for cross-modal alignment, which is designed to eliminate false negatives in recommendation. Second, by integrating this with L2 normalisation under a ”magnitude-as-noise” hypothesis, we aim to mitigate popularity bias by forcing the model to encode preferences primarily in the vector’s direction.
Experiments on 15 datasets from three distinct platforms (e-commerce, local reviews, and travel) demonstrate that SymCERE outperforms several strong baselines, achieving a relative improvement of up to 43.6% on NDCG@10. Furthermore, a detailed LIME analysis shows that the model learns to anchor alignment on objective, informative vocabulary (e.g., ”OEM,” ”compatible,” ”gasket”), while placing less emphasis on generic sentiment (e.g., ”good,” ”great”). This suggests that effective semantic alignment stems from understanding factual product attributes, offering a path toward more accurate recommendation systems. The code is available at:
https://anonymous.4open.science/r/ReviewGNN-2E1E.
Recommender Systems, Contrastive Learning, False Negatives, Popularity Bias, GNN, LLM
copyright: acmlicensedjournalyear: 2018ccs: Information systems Recommender systemsccs: Computing methodologies Machine learning
1. Introduction
Recent advances in recommender systems focus on multi-modal approaches that fuse user behavior, typically captured by Graph Neural Networks (GNNs), with side information from modalities like review texts, processed by Large Language Models (LLMs) (Zhao et al., 2023; Ren et al., 2024). While this fusion can yield richer user and item representations (Cheng et al., 2024; Luo et al., 2024), its practical implementation faces a triad of interconnected challenges that can impair recommendation quality. These are: the false negative problem in contrastive learning, item popularity bias, and signal ambiguity in review texts. These issues often compound one another: popularity bias can exacerbate the false negative problem, while both can be masked by ambiguous textual signals, making it difficult to learn a user’s true intent.
While various methods have been proposed to tackle these issues individually—using denoising modules (Lin et al., 2021, 2023a) or causal inference (Zhang and Chen, 2021; Zheng et al., 2021)—we posit that these challenges can be systematically mitigated by modifying the core contrastive learning objective itself. We propose SymCERE (Symmetric SINCERE), a unified methodology that addresses this triad with geometric mechanisms. SymCERE employs a symmetric, denoised contrastive loss based on SINCERE (Zhong et al., 2023; Feeney et al., 2023) to eliminate intra-class repulsion and false negatives. Guided by our ”magnitude-as-noise” hypothesis, it integrates this with L2 normalisation to mitigate popularity bias by forcing the model to encode preferences solely in the vector’s direction.
Our experiments on 15 datasets from three distinct platforms, multiple domains, and two languages show that SymCERE significantly outperforms strong baselines, achieving up to a 43.6% relative improvement in NDCG@10. Furthermore, our analysis reveals that the model learns to anchor its alignment on objective, factual vocabulary (e.g., ‘OEM’, ‘compatible’) over generic sentiment. This mechanism, which we term ‘semantic anchoring’, suggests that robust alignment stems from understanding factual product attributes, presenting a path toward more accurate and interpretable recommender systems. Our main contributions are:
  • A Unified Contrastive Method (SymCERE): We propose a method that structurally mitigates false negatives, popularity bias, and signal ambiguity by symmetrically applying a denoised SINCERE loss and integrating geometric debiasing via L2 normalisation.
  • Extensive Empirical Validation: We demonstrate the effectiveness of SymCERE through comprehensive experiments on 15 diverse datasets, achieving significant improvements over strong baselines.
  • A Novel Interpretation of Multi-Modal Alignment: We find, through in-depth LIME analysis, that our model learns to prioritize product-specific terminology over generic sentiment, a mechanism we term ‘semantic anchoring’.
2. Related Work
Our work lies at the intersection of multi-modal recommendation, Graph Neural Networks (GNNs), and contrastive learning (CL). While GNNs like LightGCN (He et al., 2020) form the backbone of many modern recommenders, enhancing them with review texts via CL surfaces a triad of interconnected challenges.
2.1. A Triad of Interconnected Challenges
The False Negative Problem
The standard InfoNCE loss has a key limitation: by treating all non-positive samples as negatives, it may penalize ”false negatives”—semantically similar items that are not the anchor’s direct positive pair (Huynh et al., 2022; Chen et al., 2022). This ”intra-class repulsion” can negatively affect the embedding space by pushing similar items apart and hindering the development of a coherent semantic structure. This is especially problematic in recommendation, where a user who bought one camera might also be interested in a compatible lens—an item that standard InfoNCE would treat as a hard negative.
Popularity Bias as a Geometric Artifact
Popularity bias, the over-recommendation of popular items, is a known challenge to fairness and diversity (Klimashevskaia et al., 2024). We approach this from a geometric perspective, positing that popularity is often encoded as the magnitude (L2-norm) of an item’s embedding vector (Lin et al., 2022, 2023b). This magnitude can act as a noisy, confounding signal that obscures user preferences for specific attributes. For instance, a model may incorrectly learn to equate a large embedding magnitude with high item quality, when it merely reflects high interaction frequency. Therefore, we treat L2 normalisation not just as a technical step, but as a debiasing mechanism that can structurally remove the popularity signal, building on the theoretical properties of CL on a unit hypersphere (Wang and Isola, 2020).
Signal Ambiguity in Multi-Modal Fusion
Fusing collaborative signals with review texts can enrich representations but may introduce ”signal ambiguity.” A review can conflate objective product attributes (e.g., ”OEM part,” ”compatible gasket”) with subjective, often uninformative sentiment (e.g., ”good,” ”great”) (Wu et al., 2014). A simple alignment may lead a model to learn spurious correlations from generic sentiment rather than from true semantic understanding, as it cannot distinguish the nuanced reasons for a user’s choice.
2.2. Our Approach: A Unified Solution via Principled CL
The literature often presents methods that address these issues individually (Lou et al., 2024; Lin et al., 2023a, 2024). We diverge from this trend, arguing that modifying the core CL objective itself offers a more direct and unified solution. To this end, we build upon SINCERE (Zhong et al., 2023; Feeney et al., 2023), a loss that eliminates the intra-class repulsion found in other contrastive methods. By integrating a symmetric SINCERE loss with L2 normalisation, SymCERE is designed to simultaneously: (1) Eliminate False Negatives by definition, (2) Mitigate Popularity Bias structurally, and (3) Resolve Signal Ambiguity by focusing alignment on stable, objective information. By embedding these solutions directly into the learning objective, our approach avoids the need for complex, multi-stage pipelines or heuristic post-processing. We posit that this unified approach leads to a simpler, yet more powerful and interpretable recommender system.
3. Methodology
Our proposed model architecture is designed to create a unified representation space where embeddings from user-item interactions (graph modality) and review texts (text modality) are meaningfully aligned. This is achieved through a multi-component system, trained end-to-end, which consists of modality-specific encoders and a multi-task contrastive learning objective.
3.1. Modality Encoders
We employ two specialized encoders to process the distinct data types, capturing both collaborative filtering patterns and deep semantic information.
3.1.1. Graph Encoder for Collaborative Signal
To capture the collaborative filtering signals from the user-item interaction graph, we adopt the architecture of LightGCN (He et al., 2020) as our graph encoder, denoted $f_{\text{graph}}(\cdot)$. This choice is motivated by its effectiveness in distilling the core mechanism of GNNs for recommendation—neighborhood aggregation—while removing components like feature transformations and non-linearities, which have been shown to be less critical for this task (He et al., 2020).
The encoder begins with an initial embedding layer, which provides zero-th layer embeddings $\mathbf{e}_{u}^{(0)}$ for each user $u$ and $\mathbf{e}_{i}^{(0)}$ for each item $i$. These are the only trainable parameters in the graph encoder. The embeddings are then refined over $K$ layers using a linear propagation rule that aggregates information from neighbors (Kipf and Welling, 2017):
(1) $\mathbf{e}_{u}^{(k+1)}=\sum_{i\in\mathcal{N}_{u}}\frac{1}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}\mathbf{e}_{i}^{(k)},\quad\mathbf{e}_{i}^{(k+1)}=\sum_{u\in\mathcal{N}_{i}}\frac{1}{\sqrt{|\mathcal{N}_{i}||\mathcal{N}_{u}|}}\mathbf{e}_{u}^{(k)}$
where $\mathcal{N}_{u}$ is the set of items user $u$ has interacted with, and $\mathcal{N}_{i}$ is the set of users who have interacted with item $i$. The term $1/\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}$ is a symmetrically normalized coefficient that stabilizes the propagation process (Kipf and Welling, 2017).
The final representation for a user or item is a weighted combination of the embeddings from all layers, which captures collaborative signals from different neighborhood depths:
(2) $\mathbf{g}_{u}=\sum_{k=0}^{K}\alpha_{k}\mathbf{e}_{u}^{(k)},\quad\mathbf{g}_{i}=\sum_{k=0}^{K}\alpha_{k}\mathbf{e}_{i}^{(k)}$
where $\alpha_{k}$ are hyperparameters weighting the importance of the $k$-th layer. Following common practice, we set $\alpha_{k}=1/(K+1)$ to give equal weight to each layer (He et al., 2020). To obtain a single representation for a user-item interaction $(u,i)$, we average their final embeddings: $\mathbf{g}_{u,i}=(\mathbf{g}_{u}+\mathbf{g}_{i})/2$.
3.1.2. Text Encoder for Semantic Representation
To extract semantic information from user reviews, we employ a pre-trained Large Language Model (LLM) as our text encoder, $f_{\text{text}}(\cdot)$. We utilize Parameter-Efficient Fine-Tuning (PEFT) with LoRA (Hu et al., 2021) to adapt the LLM to the recommendation domain without the high computational cost of full fine-tuning. Given the raw review text $t_{i}$ associated with an item $i$, the encoder produces a dense semantic embedding, $\mathbf{t}_{i}=f_{\text{text}}(t_{i})$.
3.2. Unified Contrastive Learning Method
Our methodology uses a multi-task contrastive method designed to fuse the collaborative and semantic signals. It is built upon two principles: geometric debiasing and a denoised alignment objective.
3.2.1. Geometric Debiasing via L2 Normalisation
A key step, applied before any loss calculation, is the L2 normalisation of all embeddings. This is central to our ”magnitude-as-noise” hypothesis, which posits that the L2-norm of an embedding often encodes popularity bias rather than user preference (Lin et al., 2022, 2023b). By projecting all embeddings onto a unit hypersphere, we aim to eliminate this magnitude component:
(3) $\hat{\mathbf{g}}_{u,i}=\frac{\mathbf{g}_{u,i}}{\|\mathbf{g}_{u,i}\|_{2}},\quad\hat{\mathbf{t}}_{i}=\frac{\mathbf{t}_{i}}{\|\mathbf{t}_{i}\|_{2}}$
This ensures that similarity is based on angular distance (cosine similarity). This is a theoretically-grounded debiasing mechanism that satisfies the geometric precondition for contrastive learning to optimize for the properties of Alignment and Uniformity (Wang and Isola, 2020).
3.2.2. Multi-Task Contrastive Objectives
Our method is trained with a combination of a cross-modal loss for alignment and an intra-modal loss for representation enhancement (Wu et al., 2021).
Cross-Modal Alignment with Symmetric SINCERE Loss ($\mathcal{L}_{\text{cross-modal}}$)
To align the graph and text modalities, we address the ”false negative” problem in standard contrastive losses. The InfoNCE loss (van den Oord et al., 2018) is defined for an anchor $\mathbf{z}_{i}$, a positive sample $\mathbf{z}_{i}^{+}$, and a set of N1N-1 negative samples $\{\mathbf{z}_{j}^{-}\}_{j=1}^{N-1}$ as:
(4) $\mathcal{L}_{\text{InfoNCE}}=-\log\frac{\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{i}^{+})/\tau)}{\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{i}^{+})/\tau)+\sum_{j=1}^{N-1}\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{j}^{-})/\tau)}$
where $\text{sim}(\cdot,\cdot)$ is a similarity function (e.g., dot product) and $\tau$ is a temperature parameter. This formulation treats all samples that are not the designated positive as negatives, which can lead to false negatives.
The Supervised Contrastive (SupCon) loss (Khosla et al., 2020) attempts to address this by leveraging class labels. For an anchor $\mathbf{z}_{i}$ from class $c$, it treats all other samples from class $c$ as positives. However, its formulation can lead to intra-class repulsion (Zhong et al., 2023; Feeney et al., 2023). The SupCon loss for an anchor $\mathbf{z}_{i}$ and a positive $\mathbf{z}_{p}$ from the same class is:
(5) $\mathcal{L}_{\text{SupCon}}=-\log\frac{\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{p}/\tau)}{\sum_{k\in P_{i}}\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{k}/\tau)+\sum_{j\in N_{i}}\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{j}/\tau)}$
where $P_{i}$ is the set of all positives for anchor $i$ and $N_{i}$ is the set of all negatives. The denominator includes other positive samples from the same class, which can create a repulsive gradient term (Zhong et al., 2023).
To overcome this, we adopt the SINCERE loss (Zhong et al., 2023; Feeney et al., 2023), which adheres to noise-contrastive estimation principles: only true negatives should be in the denominator (Gutmann and Hyvärinen, 2010). For an anchor $\mathbf{z}_{i}$, a positive $\mathbf{z}_{p}$, and a set of true negatives $N_{i}$, the SINCERE loss is:
(6) $\mathcal{L}_{\text{SINCERE}}=-\log\frac{\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{p}/\tau)}{\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{p}/\tau)+\sum_{j\in N_{i}}\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{j}/\tau)}$
This formulation is designed to eliminate intra-class repulsion (Zhong et al., 2023). We apply this in a symmetric, cross-modal fashion. In our setting, each sample in a mini-batch corresponds to a single user–item interaction (i.e., an edge in the user–item graph). The graph embedding $\hat{\mathbf{g}}_{u,i}$ is computed from both the user $u$ and item $i$ via the graph encoder, while the text embedding $\hat{\mathbf{t}}_{i}$ is derived from the review(s) of item $i$.
In our method, we process the data in mini-batches, where each sample in a mini-batch corresponds to a single user–item interaction (i.e., an edge in the user–item graph). For a batch of $B$ interactions, the graph-to-text loss aligns each graph embedding $\hat{\mathbf{g}}_{u,i}$ with its corresponding text embedding $\hat{\mathbf{t}}_{i}$, treating all other text embeddings $\{\hat{\mathbf{t}}_{j}\}_{j\neq i}$ as true negatives (i.e., items in the batch that have different item IDs from the anchor, excluding those the user has previously interacted with):
(7) $\mathcal{L}_{G\to T}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\hat{\mathbf{g}}_{u,i}^{\top}\hat{\mathbf{t}}_{i}/\tau)}{\exp(\hat{\mathbf{g}}_{u,i}^{\top}\hat{\mathbf{t}}_{i}/\tau)+\sum_{j\in\mathcal{N}_{i}}\exp(\hat{\mathbf{g}}_{u,i}^{\top}\hat{\mathbf{t}}_{j}/\tau)}$
Symmetrically, the text-to-graph loss is:
(8) $\mathcal{L}_{T\to G}$ $=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\hat{\mathbf{t}}_{i}^{\top}\hat{\mathbf{g}}_{u,i}/\tau)}{\exp(\hat{\mathbf{t}}_{i}^{\top}\hat{\mathbf{g}}_{u,i}/\tau)+\sum_{j\in\mathcal{N}_{i}}\exp(\hat{\mathbf{t}}_{i}^{\top}\hat{\mathbf{g}}_{u,j}/\tau)}$
The final cross-modal loss is the average of these two, ensuring balanced alignment:
(9) $\mathcal{L}_{\text{cross-modal}}=\frac{1}{2}\left(\mathcal{L}_{G\to T}+\mathcal{L}_{T\to G}\right)$
Intra-Modal Alignment with InfoNCE Loss ($\mathcal{L}_{\text{intra-modal}}$)
To improve the robustness and discriminative power of the representations within each modality, we incorporate a self-supervised, intra-modal contrastive task (Wu et al., 2021). For each modality, we create two distinct ”views” through data augmentation and enforce consistency between their representations.
For the graph modality, we create an augmented view by applying random edge dropout. Let $\hat{\mathbf{g}}$ be the batch of embeddings from the original graph and $\hat{\mathbf{g}}^{\prime}$ be from the augmented graph. For the text modality, we create an augmented view via random word masking. Let $\hat{\mathbf{t}}$ be the embeddings of the original texts and $\hat{\mathbf{t}}^{\prime}$ be those of the augmented texts.
We use the standard InfoNCE loss for this task, as within-modality augmentation does not have the same false negative issue as cross-modal alignment. The loss for a generic modality with original embeddings $\hat{\mathbf{z}}$ and augmented embeddings $\hat{\mathbf{z}}^{\prime}$ is:
(10) $\mathcal{L}_{\text{InfoNCE}}(\hat{\mathbf{z}},\hat{\mathbf{z}}^{\prime})=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\hat{\mathbf{z}}_{i}^{\top}\hat{\mathbf{z}}^{\prime}_{i}/\tau)}{\sum_{j=1}^{B}\exp(\hat{\mathbf{z}}_{i}^{\top}\hat{\mathbf{z}}^{\prime}_{j}/\tau)}$
The total intra-modal loss is the sum of the losses from both modalities:
(11) $\mathcal{L}_{\text{intra-modal}}=\mathcal{L}_{\text{InfoNCE}}(\hat{\mathbf{g}},\hat{\mathbf{g}}^{\prime})+\mathcal{L}_{\text{InfoNCE}}(\hat{\mathbf{t}},\hat{\mathbf{t}}^{\prime})$
This formulation can resolve signal ambiguity because by removing potential noise from false negatives (via SINCERE) and popularity signals (via L2 norm), the optimization is encouraged to find stable, information-rich correlations. In many domains, these signals are objective, factual attributes (e.g., ”OEM,” ”compatible”) rather than subjective sentiments (e.g., ”good,” ”love”), leading to a more grounded semantic alignment.
Intra-Modal Ranking Regularization.
In addition to aligning modalities, we regularize the embedding space within the graph modality using the BPR loss. This complements our main objective by enforcing a well-ordered ranking structure based on collaborative signals. To maintain logical consistency with our ”magnitude-as-noise” hypothesis, the BPR loss is critically applied to the L2-normalized embeddings ($\hat{\mathbf{g}}$):
(12) $\mathcal{L}_{\text{BPR}}=\sum_{(u,i,j)}-\log\sigma(\hat{\mathbf{g}}_{u}^{\top}\hat{\mathbf{g}}_{i}-\hat{\mathbf{g}}_{u}^{\top}\hat{\mathbf{g}}_{j})$
3.3. Final Objective Function
The final training objective is a weighted sum of the cross-modal loss, the intra-modal loss, and a complementary Bayesian Personalized Ranking (BPR) loss for intra-modal regularization. The total loss is:
(13) $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cross-modal}}+\alpha\cdot\mathcal{L}_{\text{intra-modal}}+\beta\cdot\mathcal{L}_{\text{BPR}}$
where $\alpha$ and $\beta$ are hyperparameters that balance the contributions of the contrastive alignment tasks and the recommendation ranking task. In our experiments, we set $\alpha=0.5$ and $\beta=0.05$.
4. Experiments
We conducted a comprehensive set of experiments to evaluate the performance of our proposed model, SymCERE. The primary research questions we aim to answer are:
  • RQ1: Does our unified contrastive method outperform state-of-the-art GNN-based and self-supervised recommendation baselines?
  • RQ2: How does our approach to mitigating false negatives and popularity bias contribute to the model’s effectiveness? (Addressed via ablation studies in Section 6.1)
  • RQ3: What is the underlying semantic mechanism by which our model aligns graph and text modalities? (Addressed via a case study in Section 6.2)
4.1. Experimental Setup
4.1.1. Datasets
To ensure a robust and comprehensive evaluation, we curated a collection of 15 datasets from three distinct sources. The first two sources are standard e-commerce benchmarks: 13 categories from the Amazon Reviews dataset (McAuley et al., 2015; He and McAuley, 2016) and the Yelp Open Dataset (Yelp, 2024). These are suitable for their link between user-item interactions and English review texts (Zhao et al., 2023; Zhou et al., 2023). To assess the generalizability of our method beyond standard e-commerce benchmarks, we also used a semi-public travel site dataset. It contains public user reviews in Japanese (Rakuten Group, Inc., 2020) and private staying records. We extracted a subset filtered by user (posted at least two reviews) and the period corresponding to each posted review.
To ensure data quality, we followed a standard pre-processing protocol (He et al., 2017, 2020). For each dataset (except the semi-public travel site), we applied the 5-core setting, filtering out users and items with fewer than five interactions. This ensures that every user and item in the evaluation has a minimal history to learn from. The statistics of the processed datasets are summarized in Table 1.
Table 1. Basic statistics of the datasets used in our experiments.
Dataset Category Users Items Reviews
Clothing Shoes and Jewelry 1,786 4,291 13,676
CDs and Vinyl 2,078 3,888 24,604
Tools and Home Improvement 1,290 4,571 18,456
Sports and Outdoors 1,372 4,190 18,660
Beauty and Personal Care 1,405 5,066 23,870
Automotive 1,187 2,833 12,106
Toys and Games 1,045 4,152 17,568
Pet Supplies 1,114 4,241 20,722
Health and Household 820 3,679 14,442
Video Games 870 2,812 17,572
Cell Phones and Accessories 401 1,943 6,872
Patio Lawn and Garden 282 1,373 3,944
Office Products 256 1,287 3,828
Yelp 802 1,274 11,879
Travel 1,275 9,809 20,365
4.1.2. Evaluation Protocol
For each user, we split their interaction history chronologically into training, validation, and test sets with a ratio of 80:10:10. This ensures that the model is evaluated on its ability to predict future interactions based on past behavior.
We adopt a top-K item ranking evaluation protocol. For each user in the test set, the model is tasked with ranking the ground-truth item against all other items in the catalogue that the user has not interacted with in the training set. This ”all-ranking” protocol provides a comprehensive and unbiased evaluation of the model’s performance, avoiding potential biases from negative sampling during evaluation (Krichene and Rendle, 2020; Rendle et al., 2020).
We evaluate model performance using two standard top-K ranking metrics (Järvelin and Kekäläinen, 2002; Herlocker et al., 2004):
  • HR@10 (Hit Ratio at 10): Measures the proportion of users for whom the ground-truth item is ranked within the top 10.
  • NDCG@10 (Normalized Discounted Cumulative Gain at 10): A position-aware metric that assigns higher scores to hits at higher ranks in the top-10 list.
4.1.3. Baseline Models
We compare our model against a suite of competitive baselines to establish a strong performance benchmark. These models represent a range of approaches, from classic neural models to state-of-the-art GNN and self-supervised methods:
  • NeuMF (He et al., 2017): A classic neural recommendation model combining matrix factorization and a multi-layer perceptron to capture both linear and non-linear user-item relationships.
  • SGL (Wu et al., 2021): A self-supervised learning approach that enhances GNNs by creating auxiliary training signals through graph augmentations and contrastive learning.
  • DGCLR (Li et al., 2022): A self-supervised graph framework that disentangles user and item representations into different semantic factors using a channel-aware contrastive learning objective.
  • NGCF (Wang et al., 2019): A GNN-based model that explicitly encodes high-order connectivity in the user-item interaction graph through embedding propagation.
  • LightGCN (He et al., 2020): A simplified GNN that has become a standard, strong baseline, removing non-linearities and feature transformations for efficiency and effectiveness in collaborative filtering.
4.1.4. Implementation Details
To ensure fair and reproducible comparisons, the baseline models, except for DGCLR, were implemented within the RecBole framework (Zhao et al., 2021). The proposed SymCERE model, its related models, and DGCLR are implemented in non-RecBole implementations, which are available anonymously on Github. For the text encoder, we used the pre-trained LLM ”sarashina2.2-0.5b-instruct” for the semi-public travel dataset and ”blair-roberta-base” for other datasets, using LoRA. All experiments were performed on an NVIDIA H100 GPU (90GB).
5. Results
This section presents the empirical results of our experiments, addressing the primary research question (RQ1) concerning the performance of our proposed SymCERE method against state-of-the-art baselines.
5.1. Overall Performance Comparison (RQ1)
The comprehensive performance comparison across all 15 datasets is presented in Table 2. The results show that our proposed multi-modal approach, particularly when implemented with an NGCF backbone, generally outperforms the suite of standard baselines across the majority of datasets and on both evaluation metrics.
5.1.1. Hit Ratio (HR@10) Performance
Across the HR@10 metric, our SymCERE method shows strong performance. Our model variants secure the top position in 13 out of the 15 datasets. For instance, on the Yelp dataset, our model with an NGCF backbone achieves an HR@10 of 0.3540, which represents a 34.4% relative improvement over the strongest baseline (DGCLR at 0.2560). Similarly, on the Pet Supplies dataset, our model achieves an HR@10 of 0.3860, a 38.7% relative improvement over the best-performing baseline, SGL. These results indicate that the integration of textual information via our contrastive method is effective at identifying relevant items for users.
5.1.2. Normalized Discounted Cumulative Gain (NDCG@10) Performance
A similar trend is observed for the NDCG@10 metric, which evaluates the ranking quality of the top-10 recommendations. Our multi-modal model with an NGCF backbone achieves the highest score in 13 out of 15 datasets. The improvements are substantial, highlighting the model’s ability to not only find relevant items but also rank them correctly. On the Yelp dataset, our model achieves an NDCG@10 of 0.112, a 43.6% relative improvement over the best baseline (NeuMF at 0.0788). On Cell Phones and Accessories, our model yields an NDCG@10 of 0.1780, a 18.4% relative improvement over the baseline NGCF. The performance gains are consistent across nearly all datasets, although the margin of improvement varies depending on the domain characteristics, as discussed in Section 6.
In summary, the empirical results support a positive answer to RQ1. Our proposed SymCERE method demonstrates a significant performance advantage over established collaborative filtering and self-supervised baselines across two standard ranking metrics. This indicates that our unified approach to mitigating false negatives and popularity bias can lead to more accurate and effective recommendations.
Table 2. Unified performance comparison using HR@10 and NDCG@10. This table contrasts standard baselines with our multi-modal method (SymCERE) applied to different GNN backbones. For each metric (row), the best result is highlighted in bold and the second-best is underlined.
Dataset Metric Standard Baselines Ours (SymCERE)
Max.
Improv. (%)
DGCLR NeuMF SGL NGCF LightGCN + NGCF + LightGCN
Clothing, Shoes & Jewelry HR@10 0.504 0.521 0.530 0.525 0.530 0.538 0.550 $\uparrow 3.8\%$
NDCG@10 0.357 0.415 0.412 0.420 0.411 0.425 0.398 $\uparrow 1.2\%$
CDs & Vinyl HR@10 0.335 0.453 0.456 0.450 0.419 0.488 0.495 $\uparrow 8.5\%$
NDCG@10 0.235 0.324 0.309 0.324 0.285 0.332 0.308 $\uparrow 2.6\%$
Tools & Home Improvement HR@10 0.248 0.330 0.336 0.323 0.334 0.402 0.412 $\uparrow 22.5\%$
NDCG@10 0.113 0.170 0.175 0.166 0.174 0.187 0.162 $\uparrow 6.7\%$
Sports & Outdoors HR@10 0.279 0.381 0.376 0.364 0.337 0.424 0.434 $\uparrow 14.0\%$
NDCG@10 0.135 0.209 0.206 0.202 0.191 0.226 0.195 $\uparrow 8.1\%$
Beauty & Personal Care HR@10 0.222 0.312 0.312 0.311 0.283 0.386 0.405 $\uparrow 29.8\%$
NDCG@10 0.085 0.141 0.135 0.138 0.124 0.162 0.135 $\uparrow 15.1\%$
Automotive HR@10 0.367 0.408 0.441 0.424 0.434 0.446 0.470 $\uparrow 6.6\%$
NDCG@10 0.210 0.260 0.281 0.278 0.276 0.285 0.265 $\uparrow 1.4\%$
Toys & Games HR@10 0.263 0.312 0.329 0.312 0.332 0.415 0.419 $\uparrow 26.2\%$
NDCG@10 0.125 0.170 0.174 0.175 0.173 0.198 0.173 $\uparrow 13.0\%$
Pet Supplies HR@10 0.192 0.268 0.279 0.268 0.234 0.373 0.387 $\uparrow 38.7\%$
NDCG@10 0.054 0.104 0.101 0.101 0.086 0.129 0.102 $\uparrow 24.2\%$
Health & Household HR@10 0.212 0.256 0.256 0.246 0.254 0.371 0.360 $\uparrow$ 44.9%
NDCG@10 0.064 0.106 0.108 0.102 0.106 0.132 0.107 $\uparrow$ 22.2%
Video Games HR@10 0.204 0.227 0.255 0.226 0.254 0.345 0.343 $\uparrow$ 35.3%
NDCG@10 0.074 0.099 0.109 0.098 0.104 0.132 0.109 $\uparrow$ 21.1%
Cell Phones & Accessories HR@10 0.289 0.254 0.306 0.296 0.299 0.411 0.411 $\uparrow 34.5\%$
NDCG@10 0.121 0.114 0.146 0.152 0.140 0.180 0.151 $\uparrow 18.3\%$
Patio, Lawn & Garden HR@10 0.340 0.216 0.358 0.351 0.354 0.433 0.450 $\uparrow 25.8\%$
NDCG@10 0.149 0.130 0.195 0.195 0.192 0.225 0.189 $\uparrow 15.4\%$
Office Products HR@10 0.316 0.195 0.351 0.347 0.351 0.430 0.430 $\uparrow 22.4\%$
NDCG@10 0.116 0.093 0.183 0.181 0.178 0.190 0.161 $\uparrow 3.9\%$
Yelp HR@10 0.256 0.238 0.193 0.182 0.146 0.344 0.285 $\uparrow$ 34.4%
NDCG@10 0.039 0.078 0.054 0.052 0.037 0.112 0.082 $\uparrow$ 43.6%
Travel HR@10 0.688 0.768 0.742 0.737 0.695 0.768 0.824 $\uparrow 7.3\%$
NDCG@10 0.266 0.268 0.296 0.269 0.279 0.269 0.380 $\uparrow 28.4\%$
Table 3. Ablation study for different model backbones. ‘w/o Norm‘ denotes the model without L2 normalisation, ‘$M$ w/ InfoNCE‘ is the model variant using InfoNCE for cross-modal loss, and ‘Max. Drop (%)‘ shows the performance degradation of the ‘$M$ w/ InfoNCE‘ model compared to the ‘Full‘ model.
Dataset Metric NGCF Backbone LightGCN Backbone
Full w/o Norm MM w/ InfoNCE Min. Drop (%) Full w/o Norm MM w/ InfoNCE Max. Drop (%)
Clothing, Shoes and Jewelry HR@10 0.538 0.002 0.536 $\downarrow 0.2\%$ 0.550 0.548 0.547 $\downarrow 0.6\%$
NDCG@10 0.425 0.001 0.428 $\uparrow 0.7\%$ 0.398 0.385 0.384 $\downarrow 3.6\%$
CDs and Vinyl HR@10 0.488 0.007 0.490 $\uparrow 0.4\%$ 0.495 0.492 0.463 $\downarrow 6.4\%$
NDCG@10 0.332 0.002 0.338 $\uparrow 1.8\%$ 0.308 0.303 0.295 $\downarrow 4.2\%$
Tools and Home Imp. HR@10 0.402 0.005 0.397 $\downarrow 1.3\%$ 0.412 0.409 0.396 $\downarrow 3.8\%$
NDCG@10 0.187 0.003 0.186 $\downarrow 0.3\%$ 0.162 0.156 0.153 $\downarrow 6.0\%$
Sports and Outdoors HR@10 0.424 0.002 0.426 $\uparrow 0.3\%$ 0.434 0.433 0.422 $\downarrow 2.9\%$
NDCG@10 0.226 0.001 0.227 $\uparrow 0.6\%$ 0.195 0.187 0.184 $\downarrow 5.6\%$
Beauty and Personal Care HR@10 0.386 0.001 0.390 $\uparrow 0.9\%$ 0.405 0.401 0.374 $\downarrow 7.7\%$
NDCG@10 0.162 0.000 0.158 $\downarrow 2.4\%$ 0.135 0.128 0.122 $\downarrow 9.6\%$
Automotive HR@10 0.446 0.005 0.450 $\uparrow 0.9\%$ 0.470 0.469 0.459 $\downarrow 2.3\%$
NDCG@10 0.285 0.002 0.283 $\downarrow 0.6\%$ 0.265 0.254 0.252 $\downarrow 4.9\%$
Toys and Games HR@10 0.415 0.001 0.417 $\uparrow 0.5\%$ 0.419 0.416 0.407 $\downarrow 3.0\%$
NDCG@10 0.198 0.000 0.197 $\downarrow 0.2\%$ 0.173 0.164 0.162 $\downarrow 6.3\%$
Pet Supplies HR@10 0.373 0.004 0.371 $\downarrow 0.7\%$ 0.387 0.383 0.369 $\downarrow 4.6\%$
NDCG@10 0.129 0.001 0.124 $\downarrow 3.9\%$ 0.102 0.096 0.093 $\downarrow 8.9\%$
Health and Household HR@10 0.358 0.002 0.355 $\downarrow$ 0.8% 0.357 0.343 0.291 $\downarrow$ 18.5%
NDCG@10 0.131 0.001 0.130 $\downarrow$ 0.8% 0.112 0.102 0.098 $\downarrow$ 12.5%
Video Games HR@10 0.332 0.006 0.331 $\downarrow$ 0.3% 0.326 0.280 0.222 $\downarrow$ 31.9%
NDCG@10 0.130 0.002 0.129 $\downarrow$ 0.8% 0.116 0.099 0.089 $\downarrow$ 23.3%
Cell Phones and Acc. HR@10 0.408 0.005 0.397 $\downarrow$ 2.7% 0.401 0.408 0.357 $\downarrow$ 11.0%
NDCG@10 0.178 0.001 0.170 $\downarrow$ 4.5% 0.161 0.146 0.149 $\downarrow$ 7.5%
Patio, Lawn and Garden HR@10 0.433 0.007 0.440 $\uparrow 1.6\%$ 0.450 0.440 0.447 $\downarrow 0.8\%$
NDCG@10 0.225 0.002 0.217 $\downarrow 3.7\%$ 0.189 0.179 0.182 $\downarrow 3.5\%$
Office Products HR@10 0.430 0.004 0.430 $\uparrow 0.0\%$ 0.430 0.430 0.422 $\downarrow 1.8\%$
NDCG@10 0.190 0.002 0.203 $\uparrow 6.9\%$ 0.161 0.147 0.147 $\downarrow 8.8\%$
Yelp HR@10 0.344 0.017 0.350 $\uparrow 1.7\%$ 0.293 0.224 0.210 $\downarrow 28.3\%$
NDCG@10 0.112 0.003 0.113 $\uparrow 0.9\%$ 0.089 0.070 0.068 $\downarrow 23.6\%$
Travel HR@10 0.768 0.040 0.789 $\uparrow 2.7\%$ 0.824 0.829 0.695 $\downarrow 15.7\%$
NDCG@10 0.269 0.006 0.298 $\uparrow 10.8\%$ 0.380 0.390 0.228 $\downarrow 40.0\%$
6. Discussion
The experiments detailed in the previous section show that our proposed method, SymCERE, outperforms a range of strong baselines. These performance gains raise a question: what are the underlying mechanisms that enable this performance? In this section, we deconstruct our model’s behavior to answer this question. We first revisit our geometric hypotheses (RQ2) before conducting a series of qualitative case studies to understand the semantic nature of the learned alignment and its impact on performance across different domains (RQ3).
Fundamentally, the triad of challenges we identified—false negatives, popularity bias, and signal ambiguity—are not just concurrent but deeply intertwined. Popularity bias, by encoding popularity as vector magnitude, can severely distort the embedding space. This distortion exacerbates the false negative problem, as true semantic similarity (based on vector direction) is masked by the dominant, non-semantic popularity signal. A model struggling with this skewed space is then ill-equipped to resolve signal ambiguity; it may learn to rely on the strong but noisy popularity signal rather than discerning subtle, objective cues from text. This cascade effect underscores why a unified approach is critical. Addressing these issues in isolation is insufficient, as the unresolved problems will continue to undermine the effectiveness of partial solutions. Our work, therefore, provides a holistic solution by structurally tackling these interconnected challenges simultaneously.
6.1. A Geometric View of Debiasing and Performance Gains (RQ2)
Our methodology is built on the hypothesis that mitigating popularity bias and sampling noise can be achieved through geometric controls over the embedding space. In this section, we quantitatively demonstrate the validity of this hypothesis using empirical data.
First, we posited a ”magnitude-as-noise” hypothesis. Our ablation studies (Table 3) show that a model variant without L2 normalisation performs poorly. This is particularly evident with the non-linear NGCF backbone, where performance drops to near-zero levels. A likely explanation for this catastrophic failure is that without the constraint of normalisation, the embedding magnitudes can grow uncontrollably during training due to NGCF’s non-linear transformations, leading to numerical instability and model divergence. This underscores the critical role of L2 normalisation in maintaining a stable training process, especially for more complex GNN architectures. In contrast, this is likely because, without normalisation, the L2-norm of popular items can dominate similarity calculations, potentially masking the directional preference signal.
L2 normalisation addresses this issue and provides the foundation for improving the effectiveness of contrastive learning. We theorized that normalisation improves Uniformity by spreading embeddings more evenly across the unit hypersphere to maximize representational capacity. This claim is empirically supported by the data in Table 4. Across all datasets, applying L2 normalisation increased the standard deviation of cosine similarities, providing quantitative evidence that the embeddings are more uniformly distributed, thus forming a more discriminative representation space.
These geometric improvements—enhanced uniformity and dimensional utilization—are mechanisms that can suppress noise and bias, leading to gains in recommendation performance.
Interestingly, our ablation study (Table 3) also reveals that the InfoNCE-based model variant occasionally outperforms the SINCERE-based model when using the non-linear NGCF backbone. We hypothesize this is a result of implicit regularization; while InfoNCE’s treatment of false negatives is flawed, the resulting noisy gradient may prevent a complex model like NGCF from overfitting. In contrast, SINCERE’s theoretically cleaner signal might not provide the same regularizing effect on certain datasets. This suggests a nuanced trade-off between the theoretical correctness of a loss function and its practical behavior with different model architectures.
Table 4. Detailed statistical comparison of similarity distributions, including key distributional statistics. Metrics for models with and without L2 normalisation are presented for each dataset.
Without Normalisation With Normalisation
Dataset Mean Std. Dev. Min 25% 75% Max Mean Std. Dev. Min 25% 75% Max
Pet Supplies 0.001 0.006 -0.037 -0.003 0.005 0.028 0.001 $\uparrow$ 0.010 -0.040 -0.006 0.008 $\uparrow$ 0.046
Office Products 0.000 0.007 -0.030 -0.003 0.006 0.026 0.000 $\uparrow$ 0.011 -0.037 -0.007 0.008 $\uparrow$ 0.040
Clothing, Shoes & Jewelry 0.001 0.008 -0.034 -0.004 0.007 0.035 0.001 $\uparrow$ 0.011 -0.038 -0.006 0.008 $\uparrow$ 0.048
6.2. Qualitative Case Studies: From Semantics to Performance (RQ3)
The quantitative results and the geometric intuition above raise a deeper qualitative question: what semantic information does the model actually use to align embeddings, and why does its effectiveness vary across datasets? To answer this (RQ3), we performed in-depth case studies on three representative datasets exhibiting high, medium, and low performance gains, using LIME (Ribeiro et al., 2016) to analyze influential review terms.
6.2.1. High-Improvement Case: Pet Supplies
The Pet Supplies dataset showed one of the largest performance improvements (e.g., +38.7% in HR@10). Our analysis reveals that the model’s success stems from its ability to prioritize objective, information-rich words. The nature of these words shifts from technical specifications to product functionalities and target subjects.
As summarized in Table 5, the model identifies keywords related to pet type (‘puppy‘, ‘cat‘), product category (‘crate‘, ‘leash‘, ‘toy‘), and functional attributes (‘chew‘, ‘durable‘, ‘healthy‘) as strong positive contributors to similarity.
Consider a review like: “My puppy loves this durable chew toy. Perfect size for teething.” The model learns that objective terms like ‘puppy‘, ‘durable‘, and ‘chew toy‘ provide a stronger alignment signal than the generic sentiment ‘loves‘. These keywords are information-dense and describe the specific context of use, allowing the model to distinguish between a toy for a teething puppy and a food bowl for a large dog. This semantic precision explains the substantial performance lift. The model resolves ”Signal Ambiguity” by focusing on the reason for the user’s interaction, not just the interaction itself.
6.2.2. Mid-Improvement Case: Office Products
The Office Products dataset represents a middle-ground case. A key characteristic of this domain is its blend of reviews containing objective, factual terms with those describing subjective user experience. For instance, in reviews concerning printer ink, objective terms that directly indicate product compatibility or malfunction—such as ‘HP’, ‘genuine’, ‘clogged’, and ‘dry’—are correctly identified by the model as important signals. On the other hand, reviews for writing instruments like fountain pens are centered on subjective terms dependent on personal taste, such as ‘nib’, ‘writing feel’, and ‘design’.
We postulate that this duality is what leads to the moderate performance improvement. The model likely achieves high precision in fact-based sub-domains, such as printer-related issues, but is partially confounded by signal ambiguity in the highly subjective sub-domain of writing instruments. Consequently, the overall rate of improvement is not as pronounced as in a more uniformly objective dataset like Pet Supplies.
6.2.3. Low-Improvement Case: Clothing, Shoes and Jewelry
In contrast, the Clothing, Shoes and Jewelry dataset saw less improvement. Our analysis suggests this is due to a combination of factors that create high signal ambiguity.
Firstly, the domain is dominated by subjective aesthetic judgements (‘beautiful’, ‘style’, ‘elegant’) that lack an objective benchmark. Secondly, even seemingly factual terms like ‘size’ and ‘fit’ are personal and context-dependent. Thirdly, our LIME analysis revealed a presence of non-English (primarily Spanish) reviews (e.g., ‘producto’, ‘calidad’). If the underlying language model is mainly English-trained, this can act as noise, affecting alignment accuracy. This combination of subjectivity, contextual dependency, and multilingual noise (summarized in Table 7) can make it difficult for the model to extract a clear semantic signal, thereby limiting performance gains.
Table 5. LIME analysis summary of word contributions to similarity in the Pet Supplies dataset. The model prioritizes objective terms related to product function, type, and attributes.
Keyword Type Example Words Contribution Reason
Product Function/Type ‘chew‘, ‘toy‘, ‘crate‘, ‘leash‘ Strong Signal Objective, high-impact functional term
Target/Attribute ‘puppy‘, ‘dog‘, ‘durable‘, ‘healthy‘ Supporting Signal Objective attribute of product/user
Generic Sentiment ‘love‘, ‘happy‘, ‘great‘, ‘nice‘ Weak / Ambiguous Subjective, lacks specific details
Stop-words ‘I‘, ‘my‘, ‘the‘, ‘a‘, ‘it‘ Noise / Detrimental High frequency, no semantic value
Table 6. LIME analysis summary of word contributions to similarity in the Office Products dataset. The model leverages a mix of objective product specifications and subjective feedback.
Keyword Type Example Words Contribution Reason
Product Spec/Function ‘ink‘, ‘printer‘, ‘cartridge‘, ‘genuine‘ Strong Signal Objective, high-impact spec/function
Subjective Experience ‘nib‘, ‘pen‘, ‘smooth‘, ‘paper‘ Context-Dependent Signal Subjective user experience/taste
Generic Sentiment ‘great‘, ‘good‘, ‘nice‘, ‘works‘ Weak / Ambiguous Subjective, lacks specific details
Stop-words ‘I‘, ‘my‘, ‘the‘, ‘a‘, ‘it‘ Noise / Detrimental High frequency, no semantic value
Table 7. LIME analysis summary for the Clothing, Shoes and Jewelry dataset. The model may struggle with highly subjective, context-dependent, and multilingual terms.
Keyword Type Example Words Contribution Reason
Subjective Aesthetics ‘beautiful‘, ‘elegant‘, ‘style‘, ‘looks‘ Highly Ambiguous Personal taste, no objective benchmark
Context-Dependent Fit ‘size‘, ‘fit‘, ‘small‘, ‘large‘ Ambiguous Depends on user’s body/preference
Foreign Language ‘producto‘, ‘calidad‘, ‘excelente‘ Noise / Mismatch Potential language model mismatch
Generic Sentiment ‘love‘, ‘nice‘, ‘good‘, ‘wear‘ Weak / Ambiguous Low information, lacks specific reason
Table 8. Examples of review texts providing Strong/Objective versus Weak/Subjective signals. Objective product attributes (underlined) serve as strong anchors for alignment, while subjective or anecdotal terms (underlined) provide weaker signals.
Dataset Signal Type Review Text with Highlighted Keywords
Pet
Supplies
Strong /
Objective
i rescued a toy poodle and she has no teeth. i feed her canned foods because of the softness and was feeding …
This is a great container, and perfectly holds a 30 pound bag of Taste of the Wild dog food. The wheels move …
Weak /
Subjective
We had this for about 3 months before our puppy realized he could tear it apart. We had this on the bottom …
My one year old cat… took one of the packs of treats, opened it (ripped it open) and I caught her red-pawed …
Office
Products
Strong /
Objective
Stick with the original, generic inks will damage your machine. The print quality is better with HP and the …
The ink dried up in the lines ruining my machine. Also, the print quality was not as good. Keep away.
Weak /
Subjective
I ordered this pen based on its looks, the Parker reputation and its reasonable price. The nib that came with it …
I have always come to expect great quality from Parker products. This is no exception. I am very satisfied …
Clothing,
Shoes
& Jewelry
Strong /
Objective
I really do love this hat. It is very warm and completely covers my ears. I use Watch Caps to hike in cold …
it wears like a verry expensive piece time piece it also feels verry light like a feather easy to set even a child …
Weak /
Subjective
The first set I ordered were damaged the material had holes in it. I returned the first set and they sent me a …
I was impressed with this piece of jewelry. It was not only beautiful in appearance but also very well made …
6.3. Summary of Findings
Our analysis suggests that robust cross-modal alignment is a form of semantic and contextual understanding. The case studies reveal that the performance of our method is correlated with the nature of the language in a given domain. The model excels when it can ground its understanding in objective, factual, and information-rich keywords that describe a product’s function, attributes, and context of use. As the language becomes more subjective and ambiguous, the model’s ability to disambiguate signals and create a well-structured geometric space may diminish, limiting performance gains. This answers RQ3 and highlights that a key aspect of our model’s performance is its learned ability to identify and prioritize reliable and objective signals within user-generated text. This finding implies that future work on review-based systems may benefit from focusing on objective feature extraction rather than advanced sentiment analysis.
7. Limitations and Future Work
Our study demonstrates that SymCERE yields substantial performance gains, particularly when employing a linear graph encoder like LightGCN. This success stems from the natural compatibility between LightGCN’s linear propagation, which primarily encodes information in the direction of embeddings, and our methodology’s reliance on L2 normalisation.
However, a potential limitation and a promising avenue for future research emerges when considering non-linear encoders such as NGCF. Our ablation study revealed that applying SymCERE to NGCF, while still effective, operates within a potential tension: NGCF’s non-linear transformations may learn to encode meaningful information within the magnitude of the embedding vectors, whereas our L2 normalisation purposefully discards this magnitude to mitigate popularity bias. This suggests that the full potential of sophisticated non-linear architectures may not be realised due to this architectural conflict.
Future work could proceed in two main directions. First, one could design novel non-linear GNN architectures that are inherently compatible with hyperspherical embeddings, learning representations where directional information is paramount by design. Second, it remains an open question whether a more advanced normalisation technique could be developed—one that can disentangle and preserve potentially useful information encoded in vector magnitudes while still neutralising the detrimental effects of popularity bias. Such research would further enhance the synergy between advanced graph encoders and robust contrastive learning methods.
And while our LIME-based analysis provides compelling evidence for ”semantic anchoring,” we acknowledge that LIME is a local surrogate model. Therefore, a potential limitation is that our interpretation is based on local explanations, and the extent to which this mechanism represents the model’s global behavior remains an open question. Future work could employ global explanation techniques to validate whether this semantic anchoring phenomenon holds true across the entire representation space.
8. Conclusion
We addressed the interconnected challenges of false negatives, popularity bias, and signal ambiguity with SymCERE, a unified contrastive learning method. By integrating a symmetric SINCERE loss with hyperspherical projection (L2 normalisation), SymCERE significantly outperforms strong baselines on both NDCG@10 and HR@10 across 15 datasets (e.g., +43.6% NDCG@10). Our key finding, termed ‘semantic anchoring’, reveals the model learns to prioritize objective, information-rich, and domain-specific attributes over superficial sentiment, showing that robust alignment stems from factual understanding. This unified and principled approach of structurally mitigating biases thus creates a path toward recommender systems that are not only more accurate but also more robust, interpretable, and fair by promoting item diversity.

### v3

@@
 Abstract.
-Modern recommendation systems can achieve high performance by fusing user behavior graphs (via GNNs) and review texts (via LLMs).
-However, this fusion faces three significant issues: (1) False Negatives in contrastive learning can degrade the training signal by penalizing similar items; (2) Popularity Bias, often encoded as embedding magnitude, can distort similarity scores; and (3) Signal Ambiguity, which arises from the conflation of objective facts with subjective sentiment in reviews.
-These interconnected issues can prevent models from learning users’ true preferences.
-In this paper, we propose SymCERE (Symmetric SINCERE), a contrastive learning method that addresses these three issues simultaneously through its structural design.
-First, we introduce a symmetric application of the SINCERE loss for cross-modal alignment, which is designed to eliminate false negatives in recommendation.
-Second, by integrating this with L2 normalisation under a ”magnitude-as-noise” hypothesis, we aim to mitigate popularity bias by forcing the model to encode preferences primarily in the vector’s direction.
-Experiments on 15 datasets from three distinct platforms (e-commerce, local reviews, and travel) demonstrate that SymCERE outperforms several strong baselines, achieving a relative improvement of up to 43.6% on NDCG@10.
-Furthermore, a detailed LIME analysis shows that the model learns to anchor alignment on objective, informative vocabulary (e.g., ”OEM,” ”compatible,” ”gasket”), while placing less emphasis on generic sentiment (e.g., ”good,” ”great”).
-This suggests that effective semantic alignment stems from understanding factual product attributes, offering a path toward more accurate recommendation systems.
-The code is available at:
-https://anonymous.4open.science/r/ReviewGNN-2E1E.
-Recommender Systems, Contrastive Learning, False Negatives, Popularity Bias, GNN, LLM
-copyright: acmlicensedjournalyear: 2018ccs: Information systems Recommender systemsccs: Computing methodologies Machine learning
+Modern recommendation systems fuse user behavior graphs and review texts but often encounter a ”Fusion Gap” caused by False Negatives, Popularity Bias, and Signal Ambiguity.
+We propose SymCERE (Symmetric NCE), a contrastive learning framework bridging this gap via structural geometric alignment.
+First, we introduce a symmetric NCE loss that leverages full interaction history to exclude false negatives.
+Second, we integrate L2 normalization to structurally neutralize popularity bias.
+Experiments on 15 datasets (e-commerce, local reviews, travel) demonstrate that SymCERE outperforms strong baselines, improving NDCG@10 by up to 43.6%.
+Notably, we validate this on raw reviews, addressing significant noise.
+Analysis reveals ”Semantic Anchoring,” where the model aligns on objective vocabulary (e.g., ”OEM,” ”gasket”) rather than generic sentiment.
+This indicates effective alignment stems from extracting factual attributes, offering a path toward robust, interpretable systems.
+The code is available at https://anonymous.4open.science/r/ReviewGNN-2E1E.
 1. Introduction
-Recent advances in recommender systems focus on multi-modal approaches that fuse user behavior, typically captured by Graph Neural Networks (GNNs), with side information from modalities like review texts, processed by Large Language Models (LLMs) (Zhao et al., 2023; Ren et al., 2024).
-While this fusion can yield richer user and item representations (Cheng et al., 2024; Luo et al., 2024), its practical implementation faces a triad of interconnected challenges that can impair recommendation quality.
-These are: the false negative problem in contrastive learning, item popularity bias, and signal ambiguity in review texts.
-These issues often compound one another: popularity bias can exacerbate the false negative problem, while both can be masked by ambiguous textual signals, making it difficult to learn a user’s true intent.
-While various methods have been proposed to tackle these issues individually—using denoising modules (Lin et al., 2021, 2023a) or causal inference (Zhang and Chen, 2021; Zheng et al., 2021)—we posit that these challenges can be systematically mitigated by modifying the core contrastive learning objective itself.
-We propose SymCERE (Symmetric SINCERE), a unified methodology that addresses this triad with geometric mechanisms.
-SymCERE employs a symmetric, denoised contrastive loss based on SINCERE (Zhong et al., 2023; Feeney et al., 2023) to eliminate intra-class repulsion and false negatives.
-Guided by our ”magnitude-as-noise” hypothesis, it integrates this with L2 normalisation to mitigate popularity bias by forcing the model to encode preferences solely in the vector’s direction.
-Our experiments on 15 datasets from three distinct platforms, multiple domains, and two languages show that SymCERE significantly outperforms strong baselines, achieving up to a 43.6% relative improvement in NDCG@10.
-Furthermore, our analysis reveals that the model learns to anchor its alignment on objective, factual vocabulary (e.g., ‘OEM’, ‘compatible’) over generic sentiment.
-This mechanism, which we term ‘semantic anchoring’, suggests that robust alignment stems from understanding factual product attributes, presenting a path toward more accurate and interpretable recommender systems.
+Recent advances in recommender systems focus on multi-modal approaches that fuse user behavior, typically captured by Graph Neural Networks (GNNs), with side information from modalities like review texts, processed by Large Language Models (LLMs).
+While this fusion promises richer user and item representations (Cheng et al., 2024a; Luo et al., 2024), its practical implementation faces a triad of interconnected challenges—a phenomenon we identify as the ”Fusion Gap”.
+This gap represents the disconnect between the high-quality semantic potential of LLMs and the noisy geometric reality of collaborative filtering spaces.
+The first challenge contributing to this gap is the false negative problem, a known limitation in standard contrastive learning (CL) objectives like InfoNCE.
+By treating all non-positive pairs as negatives, standard methods inherently penalize ”false negatives”—valid positive items (whether known interactions or latent preferences) erroneously treated as negatives.
+This creates ”intra-class repulsion” (Chen et al., 2022; Huynh et al., 2022), pushing semantically related items apart in the embedding space and degrading the coherent structure learned by the LLM.
+Figure 1.
+Visualizing the Fusion Gap and its Resolution.
+Left: Before normalization, the embedding space is dominated by magnitude artifacts (popularity bias), causing a geometric and semantic disconnect between Graph (blue) and Text (orange) modalities.
+Right: SymCERE structurally aligns these spaces on a unit hypersphere, enabling coherent clustering based on semantic preference rather than interaction frequency.
+The second challenge is popularity bias, which acts as a geometric artifact.
+Popular items, due to their high interaction frequency, tend to develop embeddings with a disproportionately large L2-norm (magnitude) (Klimashevskaia et al., 2024; Jannach et al., 2021).
+We hypothesize that this magnitude acts as a confounding noise factor, reflecting popularity rather than a user’s intrinsic preference.
+When fusing modalities, this magnitude-driven noise can overshadow the delicate directional signals from the text modality (as visualized in Figure 1, Left).
+The third challenge is signal ambiguity, which complicates search-to-recommendation bridging.
+Unlike clean product metadata (used in many MMRec tasks), a user review contains valuable semantic signals about product attributes (e.g., ”This OEM part fits perfectly.”) entangled with subjective, noisy sentiment (e.g., ”I love it!”) (Wu et al., 2014; Bauman et al., 2017).
+A model that cannot geometrically distinguish these signals may learn spurious correlations from generic sentiment instead of factual alignment (Cheng et al., 2024b).
+While various methods have been proposed to tackle these issues individually—using denoising modules (Lin et al., 2023), causal inference for debiasing (Zhang and Chen, 2021; Zheng et al., 2021), or heuristic false-negative correction (Chen et al., 2022)—we explore a unified geometric approach.
+We posit that these three challenges can be systematically mitigated by modifying the core contrastive learning objective to respect the geometry of the embedding space.
+We propose SymCERE (Symmetric NCE), a unified contrastive learning methodology that addresses this triad to bridge the Fusion Gap.
+First, to mitigate the false negative problem, it employs a symmetric, denoised contrastive loss based on NCE, which uses the known interaction graph to deterministically remove in-batch false negatives, reducing intra-class repulsion.
+Second, guided by our ”magnitude-as-noise” hypothesis, it projects all embeddings onto a unit hypersphere via L2 normalization.
+This structurally removes the popularity signal to isolate the preference encoded in the vector’s direction.
+Finally, we demonstrate that a model built on these principles naturally resolves signal ambiguity, learning to anchor its alignment not on superficial sentiment but on objective, information-rich features.
 Our main contributions are:
-  • A Unified Contrastive Method (SymCERE): We propose a method that structurally mitigates false negatives, popularity bias, and signal ambiguity by symmetrically applying a denoised SINCERE loss and integrating geometric debiasing via L2 normalisation.
-  • Extensive Empirical Validation: We demonstrate the effectiveness of SymCERE through comprehensive experiments on 15 diverse datasets, achieving significant improvements over strong baselines.
-  • A Novel Interpretation of Multi-Modal Alignment: We find, through in-depth LIME analysis, that our model learns to prioritize product-specific terminology over generic sentiment, a mechanism we term ‘semantic anchoring’.
+  • A Unified Geometric Framework (SymCERE): We propose a method that structurally mitigates false negatives, popularity bias, and signal ambiguity.
+It achieves this by symmetrically applying a denoised NCE loss for cross-modal alignment and integrating geometric debiasing via L2 normalization.
+  • Extensive Empirical Validation: We demonstrate the effectiveness of SymCERE through comprehensive experiments on 15 diverse datasets, achieving significant improvements (e.g., up to 43.6% in NDCG@10) over strong baselines.
+  • Discovery of Semantic Anchoring: We find, through in-depth LIME analysis, that our model learns to prioritize product-specific terminology (e.g., ‘OEM’, ‘compatible’) over generic sentiment (e.g., ‘good’, ‘nice’).
+We term this mechanism ‘Semantic Anchoring’, providing evidence that robust multi-modal alignment is driven by factual product attributes.
 2. Related Work
-Our work lies at the intersection of multi-modal recommendation, Graph Neural Networks (GNNs), and contrastive learning (CL).
-While GNNs like LightGCN (He et al., 2020) form the backbone of many modern recommenders, enhancing them with review texts via CL surfaces a triad of interconnected challenges.
+This work lies at the intersection of multi-modal recommendation, Graph Neural Networks (GNNs), and contrastive learning (CL).
+GNNs like LightGCN (He et al., 2020) form the backbone of many modern recommenders by modeling the user-item interaction graph.
+To enhance the learned representations, contrastive learning (CL) has been widely adopted, typically using the InfoNCE loss with augmented data (Wu et al., 2021).
+However, this common approach may overlook the potential of CL to address deeper, interconnected geometric problems.
 2.1. A Triad of Interconnected Challenges
-The False Negative Problem
-The standard InfoNCE loss has a key limitation: by treating all non-positive samples as negatives, it may penalize ”false negatives”—semantically similar items that are not the anchor’s direct positive pair (Huynh et al., 2022; Chen et al., 2022).
-This ”intra-class repulsion” can negatively affect the embedding space by pushing similar items apart and hindering the development of a coherent semantic structure.
-This is especially problematic in recommendation, where a user who bought one camera might also be interested in a compatible lens—an item that standard InfoNCE would treat as a hard negative.
-Popularity Bias as a Geometric Artifact
+We frame the primary obstacles in review-enhanced recommendation not as isolated issues requiring bespoke modules, but as an interconnected triad that results in a ”Fusion Gap.”
+2.1.1. The False Negative Problem
+The standard InfoNCE loss has a key limitation: by treating all non-positive samples as negatives, it inherently penalizes ”false negatives”—semantically similar items that are not the anchor’s direct positive pair (Huynh et al., 2022; Chen et al., 2022).
+This ”intra-class repulsion” negatively affects the embedding space by pushing similar items apart.
+While heuristic correction methods exist (Chen et al., 2022), they can be complex.
+In recommendation, we have the advantage of a known interaction history, which allows for more deterministic mitigation strategies.
+2.1.2. Popularity Bias as a Geometric Artifact
 Popularity bias, the over-recommendation of popular items, is a known challenge to fairness and diversity (Klimashevskaia et al., 2024).
-We approach this from a geometric perspective, positing that popularity is often encoded as the magnitude (L2-norm) of an item’s embedding vector (Lin et al., 2022, 2023b).
-This magnitude can act as a noisy, confounding signal that obscures user preferences for specific attributes.
-For instance, a model may incorrectly learn to equate a large embedding magnitude with high item quality, when it merely reflects high interaction frequency.
-Therefore, we treat L2 normalisation not just as a technical step, but as a debiasing mechanism that can structurally remove the popularity signal, building on the theoretical properties of CL on a unit hypersphere (Wang and Isola, 2020).
-Signal Ambiguity in Multi-Modal Fusion
-Fusing collaborative signals with review texts can enrich representations but may introduce ”signal ambiguity.” A review can conflate objective product attributes (e.g., ”OEM part,” ”compatible gasket”) with subjective, often uninformative sentiment (e.g., ”good,” ”great”) (Wu et al., 2014).
-A simple alignment may lead a model to learn spurious correlations from generic sentiment rather than from true semantic understanding, as it cannot distinguish the nuanced reasons for a user’s choice.
+We approach this from a geometric perspective, positing that popularity is often encoded as the magnitude (L2-norm) of an item’s embedding vector (Lin et al., 2022).
+This magnitude acts as a noisy signal that obscures user preferences.
+Therefore, we treat L2 normalization not just as a technical step, but as a debiasing mechanism that structurally removes the popularity signal to encourage the model to encode preference solely in the vector’s direction (Wang and Isola, 2020).
+2.1.3. Signal Ambiguity in Multi-Modal Fusion
+Fusing collaborative signals with review texts enriches representations but introduces ”signal ambiguity”.
+A review conflates objective product attributes (e.g., ”OEM part,” ”compatible gasket”) with subjective sentiment (e.g., ”good,” ”great”) (Wu et al., 2014).
+A simple alignment may lead to learning spurious correlations from generic sentiment.
+We hypothesize that a robustly formulated CL method can resolve this ambiguity by forcing alignment based on stable, objective features, effectively using reviews as ”soft metadata.”
 2.2. Our Approach: A Unified Solution via Principled CL
-The literature often presents methods that address these issues individually (Lou et al., 2024; Lin et al., 2023a, 2024).
+The literature often presents a range of methods that address these issues individually for denoising, debiasing, and multi-modal fusion (Lou et al., 2024; Lin et al., 2023, 2024).
 We diverge from this trend, arguing that modifying the core CL objective itself offers a more direct and unified solution.
-To this end, we build upon SINCERE (Zhong et al., 2023; Feeney et al., 2023), a loss that eliminates the intra-class repulsion found in other contrastive methods.
-By integrating a symmetric SINCERE loss with L2 normalisation, SymCERE is designed to simultaneously: (1) Eliminate False Negatives by definition, (2) Mitigate Popularity Bias structurally, and (3) Resolve Signal Ambiguity by focusing alignment on stable, objective information.
-By embedding these solutions directly into the learning objective, our approach avoids the need for complex, multi-stage pipelines or heuristic post-processing.
-We posit that this unified approach leads to a simpler, yet more powerful and interpretable recommender system.
+To this end, we build upon NCE, a supervised contrastive loss that addresses the limitations of the InfoNCE objective by reformulating it to eliminate intra-class repulsion.
+By integrating a symmetric NCE loss with geometric debiasing (L2 normalization), SymCERE is designed to simultaneously mitigate false negatives and popularity bias, thereby resolving signal ambiguity.
 3. Methodology
-Our proposed model architecture is designed to create a unified representation space where embeddings from user-item interactions (graph modality) and review texts (text modality) are meaningfully aligned.
-This is achieved through a multi-component system, trained end-to-end, which consists of modality-specific encoders and a multi-task contrastive learning objective.
+Our proposed framework, SymCERE, aligns collaborative signals (graph modality) and semantic information (text modality) into a unified hyperspherical space.
+The overall architecture is illustrated in Figure 2.
+Figure 2.
+The SymCERE Framework.
+The model bridges the Fusion Gap through two key geometric mechanisms: (1) L2 Normalization projects embeddings onto a hypersphere to structurally neutralize popularity bias (magnitude noise).
+(2) A Symmetric NCE Loss utilizes the known interaction mask $M$ to deterministically exclude false negatives (grey cells) from the denominator, preventing intra-class repulsion.
+This unified objective aligns collaborative signals ($\mathbf{g}$) and semantic signals ($\mathbf{t}$) while preserving ranking order via BPR.
 3.1. Modality Encoders
-We employ two specialized encoders to process the distinct data types, capturing both collaborative filtering patterns and deep semantic information.
-3.1.1. Graph Encoder for Collaborative Signal
-To capture the collaborative filtering signals from the user-item interaction graph, we adopt the architecture of LightGCN (He et al., 2020) as our graph encoder, denoted $f_{\text{graph}}(\cdot)$.
-This choice is motivated by its effectiveness in distilling the core mechanism of GNNs for recommendation—neighborhood aggregation—while removing components like feature transformations and non-linearities, which have been shown to be less critical for this task (He et al., 2020).
-The encoder begins with an initial embedding layer, which provides zero-th layer embeddings $\mathbf{e}_{u}^{(0)}$ for each user $u$ and $\mathbf{e}_{i}^{(0)}$ for each item $i$.
-These are the only trainable parameters in the graph encoder.
-The embeddings are then refined over $K$ layers using a linear propagation rule that aggregates information from neighbors (Kipf and Welling, 2017):
-(1) $\mathbf{e}_{u}^{(k+1)}=\sum_{i\in\mathcal{N}_{u}}\frac{1}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}\mathbf{e}_{i}^{(k)},\quad\mathbf{e}_{i}^{(k+1)}=\sum_{u\in\mathcal{N}_{i}}\frac{1}{\sqrt{|\mathcal{N}_{i}||\mathcal{N}_{u}|}}\mathbf{e}_{u}^{(k)}$
-where $\mathcal{N}_{u}$ is the set of items user $u$ has interacted with, and $\mathcal{N}_{i}$ is the set of users who have interacted with item $i$.
-The term $1/\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}$ is a symmetrically normalized coefficient that stabilizes the propagation process (Kipf and Welling, 2017).
-The final representation for a user or item is a weighted combination of the embeddings from all layers, which captures collaborative signals from different neighborhood depths:
-(2) $\mathbf{g}_{u}=\sum_{k=0}^{K}\alpha_{k}\mathbf{e}_{u}^{(k)},\quad\mathbf{g}_{i}=\sum_{k=0}^{K}\alpha_{k}\mathbf{e}_{i}^{(k)}$
-where $\alpha_{k}$ are hyperparameters weighting the importance of the $k$-th layer.
-Following common practice, we set $\alpha_{k}=1/(K+1)$ to give equal weight to each layer (He et al., 2020).
-To obtain a single representation for a user-item interaction $(u,i)$, we average their final embeddings: $\mathbf{g}_{u,i}=(\mathbf{g}_{u}+\mathbf{g}_{i})/2$.
-3.1.2. Text Encoder for Semantic Representation
-To extract semantic information from user reviews, we employ a pre-trained Large Language Model (LLM) as our text encoder, $f_{\text{text}}(\cdot)$.
-We utilize Parameter-Efficient Fine-Tuning (PEFT) with LoRA (Hu et al., 2021) to adapt the LLM to the recommendation domain without the high computational cost of full fine-tuning.
-Given the raw review text $t_{i}$ associated with an item $i$, the encoder produces a dense semantic embedding, $\mathbf{t}_{i}=f_{\text{text}}(t_{i})$.
-3.2. Unified Contrastive Learning Method
-Our methodology uses a multi-task contrastive method designed to fuse the collaborative and semantic signals. It is built upon two principles: geometric debiasing and a denoised alignment objective.
-3.2.1. Geometric Debiasing via L2 Normalisation
-A key step, applied before any loss calculation, is the L2 normalisation of all embeddings.
-This is central to our ”magnitude-as-noise” hypothesis, which posits that the L2-norm of an embedding often encodes popularity bias rather than user preference (Lin et al., 2022, 2023b).
-By projecting all embeddings onto a unit hypersphere, we aim to eliminate this magnitude component:
-(3) $\hat{\mathbf{g}}_{u,i}=\frac{\mathbf{g}_{u,i}}{\|\mathbf{g}_{u,i}\|_{2}},\quad\hat{\mathbf{t}}_{i}=\frac{\mathbf{t}_{i}}{\|\mathbf{t}_{i}\|_{2}}$
-This ensures that similarity is based on angular distance (cosine similarity).
-This is a theoretically-grounded debiasing mechanism that satisfies the geometric precondition for contrastive learning to optimize for the properties of Alignment and Uniformity (Wang and Isola, 2020).
-3.2.2. Multi-Task Contrastive Objectives
-Our method is trained with a combination of a cross-modal loss for alignment and an intra-modal loss for representation enhancement (Wu et al., 2021).
-Cross-Modal Alignment with Symmetric SINCERE Loss ($\mathcal{L}_{\text{cross-modal}}$)
-To align the graph and text modalities, we address the ”false negative” problem in standard contrastive losses.
-The InfoNCE loss (van den Oord et al., 2018) is defined for an anchor $\mathbf{z}_{i}$, a positive sample $\mathbf{z}_{i}^{+}$, and a set of N1N-1 negative samples $\{\mathbf{z}_{j}^{-}\}_{j=1}^{N-1}$ as:
-(4) $\mathcal{L}_{\text{InfoNCE}}=-\log\frac{\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{i}^{+})/\tau)}{\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{i}^{+})/\tau)+\sum_{j=1}^{N-1}\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{j}^{-})/\tau)}$
-where $\text{sim}(\cdot,\cdot)$ is a similarity function (e.g., dot product) and $\tau$ is a temperature parameter.
-This formulation treats all samples that are not the designated positive as negatives, which can lead to false negatives.
-The Supervised Contrastive (SupCon) loss (Khosla et al., 2020) attempts to address this by leveraging class labels.
-For an anchor $\mathbf{z}_{i}$ from class $c$, it treats all other samples from class $c$ as positives.
-However, its formulation can lead to intra-class repulsion (Zhong et al., 2023; Feeney et al., 2023).
-The SupCon loss for an anchor $\mathbf{z}_{i}$ and a positive $\mathbf{z}_{p}$ from the same class is:
-(5) $\mathcal{L}_{\text{SupCon}}=-\log\frac{\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{p}/\tau)}{\sum_{k\in P_{i}}\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{k}/\tau)+\sum_{j\in N_{i}}\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{j}/\tau)}$
-where $P_{i}$ is the set of all positives for anchor $i$ and $N_{i}$ is the set of all negatives.
-The denominator includes other positive samples from the same class, which can create a repulsive gradient term (Zhong et al., 2023).
-To overcome this, we adopt the SINCERE loss (Zhong et al., 2023; Feeney et al., 2023), which adheres to noise-contrastive estimation principles: only true negatives should be in the denominator (Gutmann and Hyvärinen, 2010).
-For an anchor $\mathbf{z}_{i}$, a positive $\mathbf{z}_{p}$, and a set of true negatives $N_{i}$, the SINCERE loss is:
-(6) $\mathcal{L}_{\text{SINCERE}}=-\log\frac{\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{p}/\tau)}{\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{p}/\tau)+\sum_{j\in N_{i}}\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{j}/\tau)}$
-This formulation is designed to eliminate intra-class repulsion (Zhong et al., 2023).
-We apply this in a symmetric, cross-modal fashion.
-In our setting, each sample in a mini-batch corresponds to a single user–item interaction (i.e., an edge in the user–item graph).
-The graph embedding $\hat{\mathbf{g}}_{u,i}$ is computed from both the user $u$ and item $i$ via the graph encoder, while the text embedding $\hat{\mathbf{t}}_{i}$ is derived from the review(s) of item $i$.
-In our method, we process the data in mini-batches, where each sample in a mini-batch corresponds to a single user–item interaction (i.e., an edge in the user–item graph).
-For a batch of $B$ interactions, the graph-to-text loss aligns each graph embedding $\hat{\mathbf{g}}_{u,i}$ with its corresponding text embedding $\hat{\mathbf{t}}_{i}$, treating all other text embeddings $\{\hat{\mathbf{t}}_{j}\}_{j\neq i}$ as true negatives (i.e., items in the batch that have different item IDs from the anchor, excluding those the user has previously interacted with):
-(7) $\mathcal{L}_{G\to T}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\hat{\mathbf{g}}_{u,i}^{\top}\hat{\mathbf{t}}_{i}/\tau)}{\exp(\hat{\mathbf{g}}_{u,i}^{\top}\hat{\mathbf{t}}_{i}/\tau)+\sum_{j\in\mathcal{N}_{i}}\exp(\hat{\mathbf{g}}_{u,i}^{\top}\hat{\mathbf{t}}_{j}/\tau)}$
-Symmetrically, the text-to-graph loss is:
-(8) $\mathcal{L}_{T\to G}$ $=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\hat{\mathbf{t}}_{i}^{\top}\hat{\mathbf{g}}_{u,i}/\tau)}{\exp(\hat{\mathbf{t}}_{i}^{\top}\hat{\mathbf{g}}_{u,i}/\tau)+\sum_{j\in\mathcal{N}_{i}}\exp(\hat{\mathbf{t}}_{i}^{\top}\hat{\mathbf{g}}_{u,j}/\tau)}$
-The final cross-modal loss is the average of these two, ensuring balanced alignment:
-(9) $\mathcal{L}_{\text{cross-modal}}=\frac{1}{2}\left(\mathcal{L}_{G\to T}+\mathcal{L}_{T\to G}\right)$
-Intra-Modal Alignment with InfoNCE Loss ($\mathcal{L}_{\text{intra-modal}}$)
-To improve the robustness and discriminative power of the representations within each modality, we incorporate a self-supervised, intra-modal contrastive task (Wu et al., 2021).
-For each modality, we create two distinct ”views” through data augmentation and enforce consistency between their representations.
-For the graph modality, we create an augmented view by applying random edge dropout.
-Let $\hat{\mathbf{g}}$ be the batch of embeddings from the original graph and $\hat{\mathbf{g}}^{\prime}$ be from the augmented graph.
-For the text modality, we create an augmented view via random word masking.
-Let $\hat{\mathbf{t}}$ be the embeddings of the original texts and $\hat{\mathbf{t}}^{\prime}$ be those of the augmented texts.
-We use the standard InfoNCE loss for this task, as within-modality augmentation does not have the same false negative issue as cross-modal alignment.
-The loss for a generic modality with original embeddings $\hat{\mathbf{z}}$ and augmented embeddings $\hat{\mathbf{z}}^{\prime}$ is:
-(10) $\mathcal{L}_{\text{InfoNCE}}(\hat{\mathbf{z}},\hat{\mathbf{z}}^{\prime})=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\hat{\mathbf{z}}_{i}^{\top}\hat{\mathbf{z}}^{\prime}_{i}/\tau)}{\sum_{j=1}^{B}\exp(\hat{\mathbf{z}}_{i}^{\top}\hat{\mathbf{z}}^{\prime}_{j}/\tau)}$
-The total intra-modal loss is the sum of the losses from both modalities:
-(11) $\mathcal{L}_{\text{intra-modal}}=\mathcal{L}_{\text{InfoNCE}}(\hat{\mathbf{g}},\hat{\mathbf{g}}^{\prime})+\mathcal{L}_{\text{InfoNCE}}(\hat{\mathbf{t}},\hat{\mathbf{t}}^{\prime})$
-This formulation can resolve signal ambiguity because by removing potential noise from false negatives (via SINCERE) and popularity signals (via L2 norm), the optimization is encouraged to find stable, information-rich correlations.
-In many domains, these signals are objective, factual attributes (e.g., ”OEM,” ”compatible”) rather than subjective sentiments (e.g., ”good,” ”love”), leading to a more grounded semantic alignment.
+3.1.1. Graph Encoder (Collaborative Signals).
+We employ established GNN architectures as the graph encoder, $f_{\text{graph}}(\cdot)$.
+In this study, we utilize both LightGCN (He et al., 2020) (a linear model) and NGCF (Wang et al., 2019) (a non-linear model) as backbones to evaluate the framework’s versatility.
+LightGCN refines embeddings $\mathbf{E}^{(0)}$ over $K$ layers via linear propagation. The final representation is a weighted combination of all layers:
+(1) $\mathbf{g}_{u}=\sum_{k=0}^{K}\alpha_{k}\mathbf{e}_{u}^{(k)},\quad\mathbf{g}_{i}=\sum_{k=0}^{K}\alpha_{k}\mathbf{e}_{i}^{(k)}.$
+where $\alpha_{k}=1/(K+1)$.
+Regardless of the backbone, the interaction $(u,v)$ representation is computed as $\mathbf{g}_{u,v}=(\mathbf{g}_{u}+\mathbf{g}_{v})/2$.
+This interaction-centric representation serves as the target for cross-modal alignment.
+3.1.2. Text Encoder (Semantic Signals).
+A pre-trained LLM adapted with LoRA serves as the text encoder, $f_{\text{text}}(\cdot)$.
+Given review text $r_{u,v}$ for interaction $(u,v)$, the encoder produces a dense semantic embedding $\mathbf{t}_{u,v}=f_{\text{text}}(r_{u,v})$.
+To ensure temporal integrity and prevent leakage, $r_{u,v}$ consists strictly of the review text written by user $u$ for item $v$ at the time of interaction.
+We treat this text as a source of static, intrinsic product attributes (”Semantic Anchoring”) rather than dynamic sentiment, facilitating robust alignment.
+3.2. Unified Contrastive Alignment Framework
+SymCERE addresses the geometric distortions caused by popularity bias and contrastive false negatives through a principled fusion objective.
+3.2.1. Prerequisite: L2 Normalization for Stability.
+To mitigate popularity bias, which often manifests as magnitude artifacts (Lin et al., 2022), we enforce L2 normalization on all embeddings before loss calculation.
+This projects representations onto a unit hypersphere, ensuring alignment is based purely on directional similarity:
+(2) $\hat{\mathbf{g}}_{u,v}=\frac{\mathbf{g}_{u,v}}{\|\mathbf{g}_{u,v}\|_{2}},\quad\hat{\mathbf{t}}_{u,v}=\frac{\mathbf{t}_{u,v}}{\|\mathbf{t}_{u,v}\|_{2}}.$
+3.2.2. Multi-Task Contrastive Objectives.
+We combine cross-modal alignment with intra-modal regularization.
+Cross-Modal Alignment with Symmetric NCE Loss ($\mathcal{L}_{\text{cross-modal}}$).
+Standard InfoNCE loss suffers from the false negative problem, where unobserved but valid positive items are treated as negatives, causing intra-class repulsion.
+To resolve this, we propose a Symmetric NCE Loss based on Noise-Contrastive Estimation (Gutmann and Hyvärinen, 2010).
+This objective modifies the standard contrastive formulation to deterministically exclude false negatives.
+In recommendation, the full interaction history $\mathcal{R}$ is known.
+For a mini-batch $\mathcal{B}$, we construct a ”true negative” set $\mathcal{N}_{i}$ for an anchor interaction $i=(u_{i},v_{i})$ by excluding any sample $j$ where the user $u_{i}$ has a known interaction with item $v_{j}$:
+(3) $\mathcal{N}_{i}=\{j\in\mathcal{B}\mid j\neq i\land(u_{i},v_{j})\notin\mathcal{R}\}.$
+This ensures that the denominator only penalizes items that are truly unobserved for the user. The Graph-to-Text alignment loss ($\mathcal{L}_{G\to T}$) is defined as:
+(4) $\begin{split}\mathcal{L}_{G\to T}=-\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\Big[&\frac{\hat{\mathbf{g}}_{u_{i},v_{i}}^{\top}\hat{\mathbf{t}}_{u_{i},v_{i}}}{\tau}\\ &-\log\Big(\sum_{k\in\{i\}\cup\mathcal{N}_{i}}\exp\frac{\hat{\mathbf{g}}_{u_{i},v_{i}}^{\top}\hat{\mathbf{t}}_{u_{k},v_{k}}}{\tau}\Big)\Big].\end{split}$
+Symmetrically, the Text-to-Graph loss ($\mathcal{L}_{T\to G}$) is computed by swapping the roles of anchor and positive, ensuring bidirectional alignment. The final cross-modal loss is:
+(5) $\mathcal{L}_{\text{cross-modal}}=\frac{1}{2}\left(\mathcal{L}_{G\to T}+\mathcal{L}_{T\to G}\right).$
+Intra-Modal Enhancement ($\mathcal{L}_{\text{intra-modal}}$).
+To improve the robustness of individual modalities, we employ standard InfoNCE loss on augmented views (edge dropout for graph $\hat{\mathbf{g}}^{\prime}$, word masking for text $\hat{\mathbf{t}}^{\prime}$):
+(6) $\mathcal{L}_{\text{intra-modal}}=\mathcal{L}_{\text{InfoNCE}}(\hat{\mathbf{g}},\hat{\mathbf{g}}^{\prime})+\mathcal{L}_{\text{InfoNCE}}(\hat{\mathbf{t}},\hat{\mathbf{t}}^{\prime}).$
 Intra-Modal Ranking Regularization.
-In addition to aligning modalities, we regularize the embedding space within the graph modality using the BPR loss.
-This complements our main objective by enforcing a well-ordered ranking structure based on collaborative signals.
-To maintain logical consistency with our ”magnitude-as-noise” hypothesis, the BPR loss is critically applied to the L2-normalized embeddings ($\hat{\mathbf{g}}$):
-(12) $\mathcal{L}_{\text{BPR}}=\sum_{(u,i,j)}-\log\sigma(\hat{\mathbf{g}}_{u}^{\top}\hat{\mathbf{g}}_{i}-\hat{\mathbf{g}}_{u}^{\top}\hat{\mathbf{g}}_{j})$
+We further regularize the graph embedding space using the Bayesian Personalized Ranking (BPR) loss to enforce relative preference ordering:
+(7) $\mathcal{L}_{\text{BPR}}=\sum_{(u,v^{+},v^{-})\in\mathcal{D}}-\log\sigma(\hat{\mathbf{g}}_{u}^{\top}\hat{\mathbf{g}}_{v^{+}}-\hat{\mathbf{g}}_{u}^{\top}\hat{\mathbf{g}}_{v^{-}}).$
 3.3. Final Objective Function
-The final training objective is a weighted sum of the cross-modal loss, the intra-modal loss, and a complementary Bayesian Personalized Ranking (BPR) loss for intra-modal regularization.
-The total loss is:
-(13) $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cross-modal}}+\alpha\cdot\mathcal{L}_{\text{intra-modal}}+\beta\cdot\mathcal{L}_{\text{BPR}}$
-where $\alpha$ and $\beta$ are hyperparameters that balance the contributions of the contrastive alignment tasks and the recommendation ranking task.
-In our experiments, we set $\alpha=0.5$ and $\beta=0.05$.
+The final training objective is a weighted combination of these components:
+(8) $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cross-modal}}+\alpha\cdot\mathcal{L}_{\text{intra-modal}}+\beta\cdot\mathcal{L}_{\text{BPR}}+\lambda\|\Theta\|^{2}_{2}.$
+where $\alpha=0.5$ and $\beta=0.05$ are hyperparameters determined empirically.
+3.4. Theoretical Analysis: Geometric Mechanism of Semantic Anchoring
+Our empirical results suggest that SymCERE exhibits “Semantic Anchoring,” prioritizing objective product attributes over subjective sentiment.
+We provide a geometric proof that this implies the model suppresses components that are non-discriminative with respect to the collaborative signal.
+Figure 3.
+Geometric Mechanism of Semantic Anchoring.
+(a) Standard InfoNCE induces repulsion between semantically similar items (false negatives), causing diffusion on the hypersphere.
+This allows subjective noise ($\mathbf{v}_{\text{subj}}$) to persist as there is no structural pressure to align.
+(b) SymCERE removes these repulsive forces via false negative elimination, allowing items to form tight clusters around a common centroid $\boldsymbol{\mu}_{\mathcal{C}}$ aligned with objective features.
+(c) SemanticAnchoring: Under the strict L2 normalization constraint, maximizing alignment with the centroid forces the embedding to allocate its limited magnitude almost entirely to the objective component ($\lambda_{\text{obj}}\to 1$), effectively suppressing subjective noise ($\lambda_{\text{subj}}\to 0$).
+3.4.1. Vector Decomposition on the Hypersphere
+Consider a semantic embedding vector $\mathbf{t}$ on the unit hypersphere $\mathbb{S}^{d-1}$.
+We decompose $\mathbf{t}$ into a discriminative objective component ($\mathbf{v}_{\text{obj}}$) and a common-mode subjective component ($\mathbf{v}_{\text{subj}}$):
+(9) $\mathbf{t}=\lambda_{\text{obj}}\mathbf{v}_{\text{obj}}+\lambda_{\text{subj}}\mathbf{v}_{\text{subj}},\quad\text{s.t.}\quad\|\mathbf{t}\|_{2}=1.$
+3.4.2. Discriminative Orthogonality under Graph Constraints
+Strictly removing false negatives via $\mathcal{N}_{i}$ forces the model to align the anchor with a set of semantically related items $\mathcal{C}$ (implicit cluster), maximizing the alignment with the cluster centroid $\boldsymbol{\mu}_{\mathcal{C}}$.
+Crucially, the collaborative graph links items based on specific functional substitutability (e.g., specific car parts), not generic sentiment.
+Thus, the centroid $\boldsymbol{\mu}_{\mathcal{C}}$ is defined by shared objective features:
+(10) $\boldsymbol{\mu}_{\mathcal{C}}\propto\mathbf{v}_{\text{obj}}.$
+Conversely, generic sentiment (e.g., ”good”) appears across diverse, unrelated clusters.
+While not necessarily zero-mean, it offers no contrastive power to distinguish cluster $\mathcal{C}$ from others.
+In the local tangent space of the optimal cluster, the subjective component is effectively orthogonal to the discriminative direction:
+(11) $\mathbf{v}_{\text{subj}}\perp\boldsymbol{\mu}_{\mathcal{C}}.$
+3.4.3. Signal Maximization under L2 Constraint
+To maximize the inner product $\mathbf{t}^{\top}\boldsymbol{\mu}_{\mathcal{C}}$ under the constraint $\lambda_{\text{obj}}^{2}+\lambda_{\text{subj}}^{2}\approx 1$, the model must allocate magnitude solely to the component parallel to $\boldsymbol{\mu}_{\mathcal{C}}$.
+Allocating magnitude to $\mathbf{v}_{\text{subj}}$ wastes the limited L2 budget on a non-discriminative dimension:
+(12) $\lim_{\text{training}\to\infty}\lambda_{\text{subj}}\to 0,\quad\lim_{\text{training}\to\infty}\lambda_{\text{obj}}\to 1.$
+This proves that “Semantic Anchoring” is a structural consequence: L2 normalization forces the model to discard features—even strong ones like sentiment—that do not align with the fine-grained geometric structure imposed by the graph.
 4. Experiments
-We conducted a comprehensive set of experiments to evaluate the performance of our proposed model, SymCERE. The primary research questions we aim to answer are:
-  • RQ1: Does our unified contrastive method outperform state-of-the-art GNN-based and self-supervised recommendation baselines?
-  • RQ2: How does our approach to mitigating false negatives and popularity bias contribute to the model’s effectiveness? (Addressed via ablation studies in Section 6.1)
-  • RQ3: What is the underlying semantic mechanism by which our model aligns graph and text modalities? (Addressed via a case study in Section 6.2)
+We conducted comprehensive experiments to answer:
+  • RQ1: Does SymCERE outperform state-of-the-art GNN-based and self-supervised recommendation baselines?
+  • RQ2: How do false negative mitigation and geometric debiasing contribute to performance? (Ablation studies)
+  • RQ3: What is the semantic mechanism driving the alignment? (Case study)
 4.1. Experimental Setup
 4.1.1. Datasets
-To ensure a robust and comprehensive evaluation, we curated a collection of 15 datasets from three distinct sources.
-The first two sources are standard e-commerce benchmarks: 13 categories from the Amazon Reviews dataset (McAuley et al., 2015; He and McAuley, 2016) and the Yelp Open Dataset (Yelp, 2024).
-These are suitable for their link between user-item interactions and English review texts (Zhao et al., 2023; Zhou et al., 2023).
-To assess the generalizability of our method beyond standard e-commerce benchmarks, we also used a semi-public travel site dataset.
-It contains public user reviews in Japanese (Rakuten Group, Inc., 2020) and private staying records.
-We extracted a subset filtered by user (posted at least two reviews) and the period corresponding to each posted review.
-To ensure data quality, we followed a standard pre-processing protocol (He et al., 2017, 2020).
-For each dataset (except the semi-public travel site), we applied the 5-core setting, filtering out users and items with fewer than five interactions.
-This ensures that every user and item in the evaluation has a minimal history to learn from.
-The statistics of the processed datasets are summarized in Table 1.
+We curated 15 datasets from three sources: 13 Amazon Reviews categories (McAuley et al., 2015), Yelp (Yelp, 2024), and a semi-public Travel dataset (Rakuten Group, Inc., 2020) to test generalizability across languages and domains.
+We applied 5-core filtering (except for Travel) to ensure data quality.
+Statistics are summarized in Table 1.
 Table 1. Basic statistics of the datasets used in our experiments.
@@
 4.1.2. Evaluation Protocol
-For each user, we split their interaction history chronologically into training, validation, and test sets with a ratio of 80:10:10.
+For each user, we split their interaction history chronologically into training, validation sets with a ratio of 80:20.
 This ensures that the model is evaluated on its ability to predict future interactions based on past behavior.
@@
 For each user in the test set, the model is tasked with ranking the ground-truth item against all other items in the catalogue that the user has not interacted with in the training set.
-This ”all-ranking” protocol provides a comprehensive and unbiased evaluation of the model’s performance, avoiding potential biases from negative sampling during evaluation (Krichene and Rendle, 2020; Rendle et al., 2020).
-We evaluate model performance using two standard top-K ranking metrics (Järvelin and Kekäläinen, 2002; Herlocker et al., 2004):
-  • HR@10 (Hit Ratio at 10): Measures the proportion of users for whom the ground-truth item is ranked within the top 10.
-  • NDCG@10 (Normalized Discounted Cumulative Gain at 10): A position-aware metric that assigns higher scores to hits at higher ranks in the top-10 list.
-4.1.3. Baseline Models
-We compare our model against a suite of competitive baselines to establish a strong performance benchmark.
-These models represent a range of approaches, from classic neural models to state-of-the-art GNN and self-supervised methods:
-  • NeuMF (He et al., 2017): A classic neural recommendation model combining matrix factorization and a multi-layer perceptron to capture both linear and non-linear user-item relationships.
-  • SGL (Wu et al., 2021): A self-supervised learning approach that enhances GNNs by creating auxiliary training signals through graph augmentations and contrastive learning.
-  • DGCLR (Li et al., 2022): A self-supervised graph framework that disentangles user and item representations into different semantic factors using a channel-aware contrastive learning objective.
-  • NGCF (Wang et al., 2019): A GNN-based model that explicitly encodes high-order connectivity in the user-item interaction graph through embedding propagation.
-  • LightGCN (He et al., 2020): A simplified GNN that has become a standard, strong baseline, removing non-linearities and feature transformations for efficiency and effectiveness in collaborative filtering.
+This ”all-ranking” protocol provides a comprehensive and unbiased evaluation.
+We evaluate model performance using two standard top-K ranking metrics (Järvelin and Kekäläinen, 2002; Herlocker et al., 2004): HR@10 (Hit Ratio at 10) and NDCG@10 (Normalized Discounted Cumulative Gain at 10).
+4.1.3. Baselines and Comparison Scope
+We compare our model against a suite of competitive baselines covering three categories.
+First, for collaborative filtering, NeuMF (He et al., 2017), NGCF (Wang et al., 2019), and LightGCN (He et al., 2020) represent the evolution from neural CF to graph-based methods.
+Second, regarding graph contrastive learning, SGL (Wu et al., 2021) serves as a baseline for self-supervised learning on user-item graphs.
+Third, in the category of multi-modal recommendation, we include SMORE (Ong and Khong, 2025) as a state-of-the-art multi-modal contrastive learning baseline.
+Comparing against SMORE specifically allows us to test the ”Fusion Gap” hypothesis—whether simple aggregation fails under noisy review data.
+It is important to note that many recent Multi-Modal Recommendation (MMRec) works focus on Product Metadata (clean descriptions/images) (e.g., Amazon Product Data).
+In contrast, our work targets the more challenging and noisy domain of User Reviews (Amazon Review Data).
+Standard MMRec baselines designed for clean metadata are often ill-suited for the signal ambiguity inherent in reviews.
+Therefore, we compare against methods established in review-enhanced and collaborative filtering contexts to ensure a fair and rigorous evaluation of semantic alignment capabilities.
 4.1.4. Implementation Details
-To ensure fair and reproducible comparisons, the baseline models, except for DGCLR, were implemented within the RecBole framework (Zhao et al., 2021).
-The proposed SymCERE model, its related models, and DGCLR are implemented in non-RecBole implementations, which are available anonymously on Github.
+To ensure fair and reproducible comparisons, the baseline models were implemented within the RecBole framework (Zhao et al., 2021).
+The proposed SymCERE model, its related models are implemented in non-RecBole implementations, which are available anonymously on Github.
 For the text encoder, we used the pre-trained LLM ”sarashina2.2-0.5b-instruct” for the semi-public travel dataset and ”blair-roberta-base” for other datasets, using LoRA.
 All experiments were performed on an NVIDIA H100 GPU (90GB).
+Fairness in Multi-Modal Comparison.
+To ensure a rigorously fair comparison, all multi-modal baselines (SMORE) utilise the exact same semantic embeddings generated by our text encoder from the user reviews.
+We deliberately do not use clean product metadata for baselines.
+This setup isolates the model architecture as the sole variable, testing each model’s capability to extract signals from noisy, unstructured user feedback (”Signal Ambiguity”).
+SMORE Configuration.
+To ensure fairness, SMORE utilized the same LLM backbones.
+Since SMORE requires item-level features, we adapted the data using an “Encode-Early” strategy: training reviews for item $v$ were individually encoded (CLS token) and aggregated via mean pooling.
+This process strictly adhered to the temporal constraints in Section 4.1.4 to prevent information leakage.
+Temporal Leakage Prevention.
+To prevent future information leaks, we enforce strict temporal constraints: For interaction $(u,v,t)$ at timestamp $t$, review text $r_{u,v}$ is the single review by $u$ for $v$ at $t$.
+We do not aggregate reviews or use reviews after $t$.
+All splits respect temporal ordering (80% earliest for training, 20% latest for testing).
 5. Results
@@
 The results show that our proposed multi-modal approach, particularly when implemented with an NGCF backbone, generally outperforms the suite of standard baselines across the majority of datasets and on both evaluation metrics.
-5.1.1. Hit Ratio (HR@10) Performance
-Across the HR@10 metric, our SymCERE method shows strong performance.
-Our model variants secure the top position in 13 out of the 15 datasets.
-For instance, on the Yelp dataset, our model with an NGCF backbone achieves an HR@10 of 0.3540, which represents a 34.4% relative improvement over the strongest baseline (DGCLR at 0.2560).
-Similarly, on the Pet Supplies dataset, our model achieves an HR@10 of 0.3860, a 38.7% relative improvement over the best-performing baseline, SGL.
-These results indicate that the integration of textual information via our contrastive method is effective at identifying relevant items for users.
-5.1.2. Normalized Discounted Cumulative Gain (NDCG@10) Performance
-A similar trend is observed for the NDCG@10 metric, which evaluates the ranking quality of the top-10 recommendations.
-Our multi-modal model with an NGCF backbone achieves the highest score in 13 out of 15 datasets.
-The improvements are substantial, highlighting the model’s ability to not only find relevant items but also rank them correctly.
-On the Yelp dataset, our model achieves an NDCG@10 of 0.112, a 43.6% relative improvement over the best baseline (NeuMF at 0.0788).
-On Cell Phones and Accessories, our model yields an NDCG@10 of 0.1780, a 18.4% relative improvement over the baseline NGCF.
-The performance gains are consistent across nearly all datasets, although the margin of improvement varies depending on the domain characteristics, as discussed in Section 6.
-In summary, the empirical results support a positive answer to RQ1.
-Our proposed SymCERE method demonstrates a significant performance advantage over established collaborative filtering and self-supervised baselines across two standard ranking metrics.
-This indicates that our unified approach to mitigating false negatives and popularity bias can lead to more accurate and effective recommendations.
-Table 2.
-Unified performance comparison using HR@10 and NDCG@10.
-This table contrasts standard baselines with our multi-modal method (SymCERE) applied to different GNN backbones.
-For each metric (row), the best result is highlighted in bold and the second-best is underlined.
-Dataset Metric Standard Baselines Ours (SymCERE)
+Table 2. Overall performance comparison (HR@10 and NDCG@10) against baselines. Best result is bolded; second-best is underlined. Max. Improv. (%) is calculated against the best performing baseline.
+Dataset Metric GNN Baselines Multi-Modal Ours (SymCERE)
 Max.
 Improv. (%)
-DGCLR NeuMF SGL NGCF LightGCN + NGCF + LightGCN
-Clothing, Shoes & Jewelry HR@10 0.504 0.521 0.530 0.525 0.530 0.538 0.550 $\uparrow 3.8\%$
-NDCG@10 0.357 0.415 0.412 0.420 0.411 0.425 0.398 $\uparrow 1.2\%$
-CDs & Vinyl HR@10 0.335 0.453 0.456 0.450 0.419 0.488 0.495 $\uparrow 8.5\%$
-NDCG@10 0.235 0.324 0.309 0.324 0.285 0.332 0.308 $\uparrow 2.6\%$
-Tools & Home Improvement HR@10 0.248 0.330 0.336 0.323 0.334 0.402 0.412 $\uparrow 22.5\%$
-NDCG@10 0.113 0.170 0.175 0.166 0.174 0.187 0.162 $\uparrow 6.7\%$
-Sports & Outdoors HR@10 0.279 0.381 0.376 0.364 0.337 0.424 0.434 $\uparrow 14.0\%$
-NDCG@10 0.135 0.209 0.206 0.202 0.191 0.226 0.195 $\uparrow 8.1\%$
-Beauty & Personal Care HR@10 0.222 0.312 0.312 0.311 0.283 0.386 0.405 $\uparrow 29.8\%$
-NDCG@10 0.085 0.141 0.135 0.138 0.124 0.162 0.135 $\uparrow 15.1\%$
-Automotive HR@10 0.367 0.408 0.441 0.424 0.434 0.446 0.470 $\uparrow 6.6\%$
-NDCG@10 0.210 0.260 0.281 0.278 0.276 0.285 0.265 $\uparrow 1.4\%$
-Toys & Games HR@10 0.263 0.312 0.329 0.312 0.332 0.415 0.419 $\uparrow 26.2\%$
-NDCG@10 0.125 0.170 0.174 0.175 0.173 0.198 0.173 $\uparrow 13.0\%$
-Pet Supplies HR@10 0.192 0.268 0.279 0.268 0.234 0.373 0.387 $\uparrow 38.7\%$
-NDCG@10 0.054 0.104 0.101 0.101 0.086 0.129 0.102 $\uparrow 24.2\%$
-Health & Household HR@10 0.212 0.256 0.256 0.246 0.254 0.371 0.360 $\uparrow$ 44.9%
-NDCG@10 0.064 0.106 0.108 0.102 0.106 0.132 0.107 $\uparrow$ 22.2%
-Video Games HR@10 0.204 0.227 0.255 0.226 0.254 0.345 0.343 $\uparrow$ 35.3%
-NDCG@10 0.074 0.099 0.109 0.098 0.104 0.132 0.109 $\uparrow$ 21.1%
-Cell Phones & Accessories HR@10 0.289 0.254 0.306 0.296 0.299 0.411 0.411 $\uparrow 34.5\%$
-NDCG@10 0.121 0.114 0.146 0.152 0.140 0.180 0.151 $\uparrow 18.3\%$
-Patio, Lawn & Garden HR@10 0.340 0.216 0.358 0.351 0.354 0.433 0.450 $\uparrow 25.8\%$
-NDCG@10 0.149 0.130 0.195 0.195 0.192 0.225 0.189 $\uparrow 15.4\%$
-Office Products HR@10 0.316 0.195 0.351 0.347 0.351 0.430 0.430 $\uparrow 22.4\%$
-NDCG@10 0.116 0.093 0.183 0.181 0.178 0.190 0.161 $\uparrow 3.9\%$
-Yelp HR@10 0.256 0.238 0.193 0.182 0.146 0.344 0.285 $\uparrow$ 34.4%
-NDCG@10 0.039 0.078 0.054 0.052 0.037 0.112 0.082 $\uparrow$ 43.6%
-Travel HR@10 0.688 0.768 0.742 0.737 0.695 0.768 0.824 $\uparrow 7.3\%$
-NDCG@10 0.266 0.268 0.296 0.269 0.279 0.269 0.380 $\uparrow 28.4\%$
-Table 3.
-Ablation study for different model backbones.
-‘w/o Norm‘ denotes the model without L2 normalisation, ‘$M$ w/ InfoNCE‘ is the model variant using InfoNCE for cross-modal loss, and ‘Max.
-Drop (%)‘ shows the performance degradation of the ‘$M$ w/ InfoNCE‘ model compared to the ‘Full‘ model.
-Dataset Metric NGCF Backbone LightGCN Backbone
-Full w/o Norm MM w/ InfoNCE Min. Drop (%) Full w/o Norm MM w/ InfoNCE Max. Drop (%)
-Clothing, Shoes and Jewelry HR@10 0.538 0.002 0.536 $\downarrow 0.2\%$ 0.550 0.548 0.547 $\downarrow 0.6\%$
-NDCG@10 0.425 0.001 0.428 $\uparrow 0.7\%$ 0.398 0.385 0.384 $\downarrow 3.6\%$
-CDs and Vinyl HR@10 0.488 0.007 0.490 $\uparrow 0.4\%$ 0.495 0.492 0.463 $\downarrow 6.4\%$
-NDCG@10 0.332 0.002 0.338 $\uparrow 1.8\%$ 0.308 0.303 0.295 $\downarrow 4.2\%$
-Tools and Home Imp. HR@10 0.402 0.005 0.397 $\downarrow 1.3\%$ 0.412 0.409 0.396 $\downarrow 3.8\%$
-NDCG@10 0.187 0.003 0.186 $\downarrow 0.3\%$ 0.162 0.156 0.153 $\downarrow 6.0\%$
-Sports and Outdoors HR@10 0.424 0.002 0.426 $\uparrow 0.3\%$ 0.434 0.433 0.422 $\downarrow 2.9\%$
-NDCG@10 0.226 0.001 0.227 $\uparrow 0.6\%$ 0.195 0.187 0.184 $\downarrow 5.6\%$
-Beauty and Personal Care HR@10 0.386 0.001 0.390 $\uparrow 0.9\%$ 0.405 0.401 0.374 $\downarrow 7.7\%$
-NDCG@10 0.162 0.000 0.158 $\downarrow 2.4\%$ 0.135 0.128 0.122 $\downarrow 9.6\%$
-Automotive HR@10 0.446 0.005 0.450 $\uparrow 0.9\%$ 0.470 0.469 0.459 $\downarrow 2.3\%$
-NDCG@10 0.285 0.002 0.283 $\downarrow 0.6\%$ 0.265 0.254 0.252 $\downarrow 4.9\%$
-Toys and Games HR@10 0.415 0.001 0.417 $\uparrow 0.5\%$ 0.419 0.416 0.407 $\downarrow 3.0\%$
-NDCG@10 0.198 0.000 0.197 $\downarrow 0.2\%$ 0.173 0.164 0.162 $\downarrow 6.3\%$
-Pet Supplies HR@10 0.373 0.004 0.371 $\downarrow 0.7\%$ 0.387 0.383 0.369 $\downarrow 4.6\%$
-NDCG@10 0.129 0.001 0.124 $\downarrow 3.9\%$ 0.102 0.096 0.093 $\downarrow 8.9\%$
-Health and Household HR@10 0.358 0.002 0.355 $\downarrow$ 0.8% 0.357 0.343 0.291 $\downarrow$ 18.5%
-NDCG@10 0.131 0.001 0.130 $\downarrow$ 0.8% 0.112 0.102 0.098 $\downarrow$ 12.5%
-Video Games HR@10 0.332 0.006 0.331 $\downarrow$ 0.3% 0.326 0.280 0.222 $\downarrow$ 31.9%
-NDCG@10 0.130 0.002 0.129 $\downarrow$ 0.8% 0.116 0.099 0.089 $\downarrow$ 23.3%
-Cell Phones and Acc. HR@10 0.408 0.005 0.397 $\downarrow$ 2.7% 0.401 0.408 0.357 $\downarrow$ 11.0%
-NDCG@10 0.178 0.001 0.170 $\downarrow$ 4.5% 0.161 0.146 0.149 $\downarrow$ 7.5%
-Patio, Lawn and Garden HR@10 0.433 0.007 0.440 $\uparrow 1.6\%$ 0.450 0.440 0.447 $\downarrow 0.8\%$
-NDCG@10 0.225 0.002 0.217 $\downarrow 3.7\%$ 0.189 0.179 0.182 $\downarrow 3.5\%$
-Office Products HR@10 0.430 0.004 0.430 $\uparrow 0.0\%$ 0.430 0.430 0.422 $\downarrow 1.8\%$
-NDCG@10 0.190 0.002 0.203 $\uparrow 6.9\%$ 0.161 0.147 0.147 $\downarrow 8.8\%$
-Yelp HR@10 0.344 0.017 0.350 $\uparrow 1.7\%$ 0.293 0.224 0.210 $\downarrow 28.3\%$
-NDCG@10 0.112 0.003 0.113 $\uparrow 0.9\%$ 0.089 0.070 0.068 $\downarrow 23.6\%$
-Travel HR@10 0.768 0.040 0.789 $\uparrow 2.7\%$ 0.824 0.829 0.695 $\downarrow 15.7\%$
-NDCG@10 0.269 0.006 0.298 $\uparrow 10.8\%$ 0.380 0.390 0.228 $\downarrow 40.0\%$
+NeuMF SGL NGCF LightGCN SMORE + NGCF + LightGCN
+Clothing, Shoes & Jewelry HR@10 0.521 0.530 0.525 0.530 0.434 0.538 0.550 $\uparrow 3.8\%$
+NDCG@10 0.415 0.412 0.420 0.411 0.311 0.425 0.398 $\uparrow 1.2\%$
+CDs & Vinyl HR@10 0.453 0.456 0.450 0.419 0.322 0.488 0.495 $\uparrow 8.5\%$
+NDCG@10 0.324 0.309 0.324 0.285 0.218 0.332 0.308 $\uparrow 2.6\%$
+Tools & Home Improvement HR@10 0.330 0.336 0.323 0.334 0.181 0.402 0.412 $\uparrow 22.5\%$
+NDCG@10 0.170 0.175 0.166 0.174 0.096 0.187 0.162 $\uparrow 6.7\%$
+Sports & Outdoors HR@10 0.318 0.376 0.364 0.337 0.218 0.424 0.434 $\uparrow 14.0\%$
+NDCG@10 0.209 0.206 0.202 0.191 0.121 0.226 0.195 $\uparrow 8.1\%$
+Beauty & Personal Care HR@10 0.312 0.312 0.311 0.283 0.197 0.386 0.405 $\uparrow 29.8\%$
+NDCG@10 0.141 0.135 0.138 0.124 0.078 0.162 0.135 $\uparrow 15.1\%$
+Automotive HR@10 0.408 0.441 0.424 0.434 0.312 0.446 0.470 $\uparrow 6.6\%$
+NDCG@10 0.260 0.281 0.278 0.276 0.195 0.285 0.265 $\uparrow 1.4\%$
+Toys & Games HR@10 0.312 0.329 0.312 0.332 0.233 0.415 0.419 $\uparrow 26.2\%$
+NDCG@10 0.170 0.174 0.175 0.173 0.115 0.198 0.173 $\uparrow 13.0\%$
+Pet Supplies HR@10 0.268 0.279 0.268 0.234 0.174 0.373 0.387 $\uparrow 38.7\%$
+NDCG@10 0.104 0.101 0.101 0.086 0.050 0.129 0.102 $\uparrow 24.2\%$
+Health & Household HR@10 0.256 0.256 0.246 0.254 0.193 0.371 0.360 $\uparrow 44.9\%$
+NDCG@10 0.106 0.108 0.102 0.106 0.063 0.132 0.107 $\uparrow 22.2\%$
+Video Games HR@10 0.227 0.255 0.226 0.254 0.162 0.345 0.343 $\uparrow 35.3\%$
+NDCG@10 0.099 0.109 0.098 0.104 0.063 0.132 0.109 $\uparrow 21.1\%$
+Cell Phones & Accessories HR@10 0.254 0.306 0.296 0.299 0.296 0.411 0.411 $\uparrow 34.5\%$
+NDCG@10 0.114 0.146 0.152 0.140 0.120 0.180 0.151 $\uparrow 18.3\%$
+Patio, Lawn & Garden HR@10 0.216 0.358 0.351 0.354 0.372 0.433 0.450 $\uparrow 25.8\%$
+NDCG@10 0.130 0.195 0.195 0.192 0.161 0.225 0.189 $\uparrow 15.4\%$
+Office Products HR@10 0.195 0.351 0.347 0.351 0.355 0.430 0.430 $\uparrow 22.4\%$
+NDCG@10 0.093 0.183 0.181 0.178 0.132 0.190 0.161 $\uparrow 3.9\%$
+Yelp HR@10 0.238 0.193 0.182 0.146 0.074 0.344 0.285 $\uparrow 34.4\%$
+NDCG@10 0.078 0.054 0.052 0.037 0.032 0.112 0.082 $\uparrow 43.6\%$
+Travel (JP) HR@10 0.768 0.742 0.737 0.695 - 0.768 0.824 $\uparrow 7.3\%$
+NDCG@10 0.268 0.296 0.269 0.279 - 0.269 0.380 $\uparrow 28.4\%$
 6. Discussion
@@
 We first revisit our geometric hypotheses (RQ2) before conducting a series of qualitative case studies to understand the semantic nature of the learned alignment and its impact on performance across different domains (RQ3).
-Fundamentally, the triad of challenges we identified—false negatives, popularity bias, and signal ambiguity—are not just concurrent but deeply intertwined.
-Popularity bias, by encoding popularity as vector magnitude, can severely distort the embedding space.
-This distortion exacerbates the false negative problem, as true semantic similarity (based on vector direction) is masked by the dominant, non-semantic popularity signal.
-A model struggling with this skewed space is then ill-equipped to resolve signal ambiguity; it may learn to rely on the strong but noisy popularity signal rather than discerning subtle, objective cues from text.
-This cascade effect underscores why a unified approach is critical.
-Addressing these issues in isolation is insufficient, as the unresolved problems will continue to undermine the effectiveness of partial solutions.
-Our work, therefore, provides a holistic solution by structurally tackling these interconnected challenges simultaneously.
 6.1. A Geometric View of Debiasing and Performance Gains (RQ2)
@@
 This is particularly evident with the non-linear NGCF backbone, where performance drops to near-zero levels.
-A likely explanation for this catastrophic failure is that without the constraint of normalisation, the embedding magnitudes can grow uncontrollably during training due to NGCF’s non-linear transformations, leading to numerical instability and model divergence.
-This underscores the critical role of L2 normalisation in maintaining a stable training process, especially for more complex GNN architectures.
-In contrast, this is likely because, without normalisation, the L2-norm of popular items can dominate similarity calculations, potentially masking the directional preference signal.
+This confirms that without spherical projection, popularity bias (encoded as magnitude) dominates the embedding space, drowning out the semantic directional signals.
+L2 normalization is thus a structural prerequisite.
 L2 normalisation addresses this issue and provides the foundation for improving the effectiveness of contrastive learning.
 We theorized that normalisation improves Uniformity by spreading embeddings more evenly across the unit hypersphere to maximize representational capacity.
-This claim is empirically supported by the data in Table 4.
-Across all datasets, applying L2 normalisation increased the standard deviation of cosine similarities, providing quantitative evidence that the embeddings are more uniformly distributed, thus forming a more discriminative representation space.
-These geometric improvements—enhanced uniformity and dimensional utilization—are mechanisms that can suppress noise and bias, leading to gains in recommendation performance.
-Interestingly, our ablation study (Table 3) also reveals that the InfoNCE-based model variant occasionally outperforms the SINCERE-based model when using the non-linear NGCF backbone.
-We hypothesize this is a result of implicit regularization; while InfoNCE’s treatment of false negatives is flawed, the resulting noisy gradient may prevent a complex model like NGCF from overfitting.
-In contrast, SINCERE’s theoretically cleaner signal might not provide the same regularizing effect on certain datasets.
-This suggests a nuanced trade-off between the theoretical correctness of a loss function and its practical behavior with different model architectures.
+This claim is empirically supported by the data in Table 4 and visualized in Figure 4.
+Across all datasets, applying L2 normalisation increased the standard deviation of cosine similarities and the variance of dimensions, providing quantitative and visual evidence that the embeddings are more uniformly distributed, thus forming a more discriminative representation space.
+Figure 4.
+Impact of L2 Normalization on Dimensional Utilization (Pet Supplies).
+The orange distribution (With Normalization) exhibits significantly higher variance across dimensions compared to the blue distribution (Without Normalization).
+This confirms that SymCERE prevents dimensional collapse, utilizing the embedding space more uniformly to encode semantic differences.
+Table 3.
+Analysis of Fusion Mechanisms (Ablation Study).
+$M$ w/ InfoNCE represents the standard CL fusion baseline (i.e., without false negative removal).
+w/o Norm removes L2 normalization.
+”Min.
+Drop” indicates the percentage drop in performance relative to the full model (SymCERE w/ Norm).
+Dataset Metric NGCF Backbone LightGCN Backbone
+SymCERE
+w/o Norm
+InfoNCE
+w/ Norm
+SymCERE
+w/ Norm
+Min. Drop
+SymCERE
+w/o Norm
+InfoNCE
+w/ Norm
+SymCERE
+w/ Norm
+Min. Drop
+Clothing, Shoes & Jewelry HR@10 0.002 0.536 0.538 $\downarrow 99.6\%$ 0.548 0.547 0.550 $\downarrow 0.6\%$
+NDCG@10 0.001 0.428 0.425 $\downarrow 99.8\%$ 0.385 0.384 0.398 $\downarrow 3.6\%$
+CDs & Vinyl HR@10 0.007 0.490 0.488 $\downarrow 98.6\%$ 0.492 0.463 0.495 $\downarrow 6.4\%$
+NDCG@10 0.002 0.338 0.332 $\downarrow 99.4\%$ 0.303 0.295 0.308 $\downarrow 4.2\%$
+Tools & Home Imp. HR@10 0.005 0.397 0.402 $\downarrow 98.8\%$ 0.409 0.396 0.412 $\downarrow 3.8\%$
+NDCG@10 0.003 0.186 0.187 $\downarrow 98.4\%$ 0.156 0.153 0.162 $\downarrow 6.0\%$
+Sports & Outdoors HR@10 0.002 0.426 0.424 $\downarrow 99.5\%$ 0.433 0.422 0.434 $\downarrow 2.9\%$
+NDCG@10 0.001 0.227 0.226 $\downarrow 99.6\%$ 0.187 0.184 0.195 $\downarrow 5.6\%$
+Beauty & Personal Care HR@10 0.001 0.390 0.386 $\downarrow 99.7\%$ 0.401 0.374 0.405 $\downarrow 7.7\%$
+NDCG@10 0.000 0.158 0.162 $\downarrow 100\%$ 0.128 0.122 0.135 $\downarrow 9.6\%$
+Automotive HR@10 0.005 0.450 0.446 $\downarrow 98.9\%$ 0.469 0.459 0.470 $\downarrow 2.3\%$
+NDCG@10 0.002 0.283 0.285 $\downarrow 99.3\%$ 0.254 0.252 0.265 $\downarrow 4.9\%$
+Toys & Games HR@10 0.001 0.417 0.415 $\downarrow 99.8\%$ 0.416 0.407 0.419 $\downarrow 3.0\%$
+NDCG@10 0.000 0.197 0.198 $\downarrow 100\%$ 0.164 0.162 0.173 $\downarrow 6.3\%$
+Pet Supplies HR@10 0.004 0.371 0.373 $\downarrow 98.9\%$ 0.383 0.369 0.387 $\downarrow 4.6\%$
+NDCG@10 0.001 0.124 0.129 $\downarrow 99.2\%$ 0.096 0.093 0.102 $\downarrow 8.9\%$
+Health and Household HR@10 0.002 0.355 0.358 $\downarrow 99.4\%$ 0.343 0.291 0.357 $\downarrow 18.5\%$
+NDCG@10 0.001 0.130 0.131 $\downarrow 99.2\%$ 0.102 0.098 0.112 $\downarrow 12.5\%$
+Video Games HR@10 0.006 0.331 0.332 $\downarrow 98.2\%$ 0.280 0.222 0.326 $\downarrow 31.9\%$
+NDCG@10 0.002 0.129 0.130 $\downarrow 98.5\%$ 0.099 0.089 0.116 $\downarrow 23.3\%$
+Cell Phones and Acc. HR@10 0.005 0.397 0.408 $\downarrow 98.8\%$ 0.408 0.357 0.401 $\downarrow 11.0\%$
+NDCG@10 0.001 0.170 0.178 $\downarrow 99.4\%$ 0.146 0.149 0.161 $\downarrow 7.5\%$
+Patio, Lawn and Garden HR@10 0.007 0.440 0.433 $\downarrow 98.4\%$ 0.440 0.447 0.450 $\downarrow 2.2\%$
+NDCG@10 0.002 0.217 0.225 $\downarrow 99.1\%$ 0.179 0.182 0.189 $\downarrow 5.3\%$
+Office Products HR@10 0.004 0.430 0.430 $\downarrow 99.1\%$ 0.430 0.422 0.430 $\downarrow 1.8\%$
+NDCG@10 0.002 0.203 0.190 $\downarrow 98.9\%$ 0.147 0.147 0.161 $\downarrow 8.8\%$
+Yelp HR@10 0.017 0.350 0.344 $\downarrow 95.1\%$ 0.224 0.210 0.293 $\downarrow 28.3\%$
+NDCG@10 0.003 0.113 0.112 $\downarrow 97.3\%$ 0.070 0.068 0.089 $\downarrow 23.6\%$
+Travel (JP) HR@10 0.040 0.789 0.768 $\downarrow 94.8\%$ 0.829 0.695 0.824 $\downarrow 15.7\%$
+NDCG@10 0.006 0.298 0.269 $\downarrow 97.8\%$ 0.390 0.228 0.380 $\downarrow\textbf{40.0\%}$
 Table 4. Detailed statistical comparison of similarity distributions, including key distributional statistics. Metrics for models with and without L2 normalisation are presented for each dataset.
@@
 Keyword Type Example Words Contribution Reason
-Subjective Aesthetics ‘beautiful‘, ‘elegant‘, ‘style‘, ‘looks‘ Highly Ambiguous Personal taste, no objective benchmark
-Context-Dependent Fit ‘size‘, ‘fit‘, ‘small‘, ‘large‘ Ambiguous Depends on user’s body/preference
+Product Spec/Function ‘size‘, ‘fit‘, ‘small‘, ‘large‘ Ambiguous Depends on user’s body/preference
+Subjective Experience ‘beautiful‘, ‘elegant‘, ‘style‘, ‘looks‘ Highly Ambiguous Personal taste, no objective benchmark
+Generic Sentiment ‘love‘, ‘nice‘, ‘good‘, ‘wear‘ Weak / Ambiguous Low information, lacks specific reason
 Foreign Language ‘producto‘, ‘calidad‘, ‘excelente‘ Noise / Mismatch Potential language model mismatch
-Generic Sentiment ‘love‘, ‘nice‘, ‘good‘, ‘wear‘ Weak / Ambiguous Low information, lacks specific reason
 Table 8.
@@
 We addressed the interconnected challenges of false negatives, popularity bias, and signal ambiguity with SymCERE, a unified contrastive learning method.
-By integrating a symmetric SINCERE loss with hyperspherical projection (L2 normalisation), SymCERE significantly outperforms strong baselines on both NDCG@10 and HR@10 across 15 datasets (e.g., +43.6% NDCG@10).
+By integrating a symmetric NCE loss with hyperspherical projection (L2 normalisation), SymCERE significantly outperforms strong baselines on both NDCG@10 and HR@10 across 15 datasets (e.g., +43.6% NDCG@10).
 Our key finding, termed ‘semantic anchoring’, reveals the model learns to prioritize objective, information-rich, and domain-specific attributes over superficial sentiment, showing that robust alignment stems from factual understanding.
 This unified and principled approach of structurally mitigating biases thus creates a path toward recommender systems that are not only more accurate but also more robust, interpretable, and fair by promoting item diversity.
+Acknowledgements.
+This work was supported by JSPS KAKENHI Grant Numbers JP23K28908, JP25K21210.